### v2

# Local EGOP for Continuous Index Learning

Abstract
We introduce the setting of continuous index learning, in which a function of many variables varies only along a small number of directions at each point. For efficient estimation, it is beneficial for a learning algorithm to adapt, near each point $x$, to the subspace that captures the local variability of the function $f$. We pose this task as kernel adaptation along a manifold with noise, and introduce Local EGOP learning, a recursive algorithm that utilizes the Expected Gradient Outer Product (EGOP) quadratic form as both a metric and inverse-covariance of our target distribution. We prove that Local EGOP learning adapts to the regularity of the function of interest, showing that under a supervised noisy manifold hypothesis, intrinsic dimensional learning rates are achieved for arbitrarily high-dimensional noise. Empirically, we compare our algorithm to the feature learning capabilities of deep learning. Additionally, we demonstrate improved regression quality compared to two-layer neural networks in the continuous single-index setting.
1 Introduction
Kernel methods are a powerful mechanism for studying machine learning algorithms. Many algorithms have leveraged the corresponding RKHS structure for efficient estimation of sufficiently regular functions (Wainwright, 2019), and functionals (Rao, 2014). Further, many popular learning algorithms, such as certain neural network architectures and random forests, have been shown to asymptotically correspond to carefully chosen kernels (Jacot et al., 2018; Scornet, 2016). Thus, the process by which investigators select kernels tailored to problems of interest, known as kernel engineering, is of central importance (Belkin et al., 2018). Beyond allowing for potential efficiency gains, kernel engineering closely emulates the feature learning properties of deep neural networks.
Some objectives of kernel engineering include the design of specialized kernels suited to particular data structures ((Odone et al., 2005; Kondor and Jebara, 2003; Joachims, 1998; Chapelle et al., 1999; Barla et al., 2002; Vishwanathan et al., 2010), etc.), statistical principles ((Genton, 2001; Schölkopf et al., 1997; Osborne, 2010), etc.), and problems of interest (Gong et al. (2024); Kokot and Luedtke (2025), etc.). In regression settings, a classical approach is local feature learning, in which kernels are augmented by differential information at points of interest ((Schmid and Mohr, 1997; Wallraven et al., 2003; Lowe, 1999)). Earlier nonparametric methods developed a similar framework, with datasets being recursively partitioned to improve the quality of local fits (Heise, 1971; Breiman and Meisel, 1976; Friedman, 1979).
Figure 1: Localizations from Local EGOP Learning (Algorithm 1) centered at each of the highlighted points (in red) trained on a grayscale image of a mandrill Bush (2021). Here $X$ is the pixel location and $Y$ the grayscale value of the image. For visualization purposes the procedure was stopped early to enforce large localizations, and the 125 highest weight pixels are highlighted at each point. On the right the image is magnified to the highlighted region boxed-off on the left.
A modern incarnation of kernel engineering is multi-index learning, particularly in the case of neural networks ((Mousavi-Hosseini et al., 2022; Boix-Adsera et al., 2023; Damian et al., 2023), etc.). This literature aims to show that the desirable properties of kernel engineering, such as data adaptivity and dimension reduction, are captured implicitly by certain machine learning models. Much work has been done in the single-index case, where the goal is to regress labels $Y$ on features $X$ when $f(x):=\mathbb{E}[Y|X=x]$ depends on $x$ solely through their evaluation in a fixed direction $v,$ via the quantity $v^{T}x$. It has been shown that two-layer neural networks not only learn this dependence, but also do so efficiently, leading to rapid increases in prediction quality ((Bietti et al., 2022; Abbe et al., 2024; Lee et al., 2024), etc.). Recent results indicate that these same benefits carry over to the multi-index setting for models trained via SGD (Damian et al., 2023; Arnaboldi et al., 2024).
In this paper, we generalize multi-index learning, introducing a new setting we call continuous-index learning (CIL). Rather than encoding the dependence of $x$ through the action of a low-rank matrix $V,$ $f(x)=g(Vx)$, we allow $V_{x}$ to be a continuously varying affine map, $f(x)=g(V_{x}x)$. Our window into this setting is via the supervised noisy manifold hypothesis, where we assume that the features $X$ are concentrated about a $d$-dimensional manifold $\mathcal{M}$, and $f(x)=g(\pi(x))$ for $\pi(x):=\operatorname{argmin}_{q\in\mathcal{M}}\|q-x\|$ the nearest point projection onto $\mathcal{M}$. For such data, the normal space $\mathcal{N}_{x}$ represents a continuously varying uninformative subspace, that is, for $\eta$ normal to $\mathcal{M}$ at $p:=\pi(x)$, $\|\eta\|$ sufficiently small, $f(x+\eta)=f(x)$. The tangent $\mathcal{T}_{x}=\mathcal{N}_{x}^{c}$ we refer to as the informative subspace. This feature structure is typical in manifold learning settings (Aamari and Levrard, 2018; Genovese et al., 2012; Kokot et al., 2025). The supervised noisy manifold hypothesis appears under a different name in the recent work Wu and Maggioni (2024), where the analysis is specialized to a 1-dimensional intrinsic manifold, a setting they call the “nonlinear single-variable model”.
Figure 2: Supervised noisy manifold data in $\mathbb{R}^{2}$. In a neighborhood about a closed curve we overlay a heatmap of a function invariant to normal displacement. The curve is displayed in black, the normal spaces are highlighted in white, and red gradients are displayed at randomly sampled points, with lengths proportionate to the gradient magnitude.
We approach this problem via kernel smoothing, $\hat{f}(x)=\sum_{i}w_{i}Y_{i}$, $w_{i}\propto k(\|x-X_{i}\|)$ for a rbf $k$. By introducing a Mahalanobis metrization $k(\|M^{1/2}(x-X_{i})\|)$, we seek to emulate the local multi-index structure of a continuous-index outcome. In particular, if $M$ can be selected to degenerate along the normal then the weights will decay primarily along the informative tangent direction, leading to estimation rates scaling with the intrinsic manifold dimension $d$. Without a priori knowledge of the manifold $\mathcal{M}$, we develop a recursive procedure we call Local EGOP Learning, which iteratively refines the metric $M$ by pooling coarse differential information estimated from the observed data.
Our method bears particular resemblance to “kernel steering”, which was developed in the image processing literature (Takeda et al., 2007) (see also the follow-up paper (Takeda et al., 2008)). The goal of kernel steering is to allow the kernel size and shape to change in a data-dependent way, adapting not only to sample location and density, but also to local features in the data. A special case of these data-adaptive kernels appears in conjunction with the popular bilateral filter in computer vision (Tomasi and Manduchi, 1998), (Elad, 2002). In Takeda et al. (2007), they propose an iterative procedure that estimates gradients about a point of interest. These are then used to “steer” the kernel locally, adopting the empirical covariance of the gradients $\hat{\mathcal{L}}$ near the point of interest as a Mahalanobis distance for subsequent estimation. Applying this method to the Gaussian kernel gives the steering kernel
$K_{h,\hat{\mathcal{L}}}(x_{i}-x^{*})\propto\exp{\left\{-\frac{(x_{i}-x^{*})^{T}\hat{\mathcal{L}}(x_{i}-x^{*})}{2h^{2}\mu^{2}}\right\}},$
which can then be used to estimate the function again and further refine the estimate of $\hat{\mathcal{L}}$. To the best of our knowledge, concrete guarantees have yet to be proven for this algorithm. This directional adaptation enables more effective denoising and image recovery, and is closely related to the expected gradient outer product (EGOP) $\mathcal{L}(\mu):=\mathbb{E}_{\mu}[\nabla f(X)\nabla f(X)^{T}]$ (Samarov, 1993; Hristache et al., 2001; Xia et al., 2002; Trivedi et al., 2014; Yuan et al., 2023). Besides kernel steering, local EGOP metrizations have appeared sporadically in different literatures, with common goals being to accelerate learning tasks and learn efficient dimension reductions (Trivedi et al., 2014; Wu et al., 2010; Mukherjee et al., 2010).
Recently, a similar procedure proposed in works such as Radhakrishnan et al. (2022, 2025); Zhu et al. (2025) emphasizes the importance of the Average Gradient Outerproduct (AGOP), which is the empirical covariance matrix of the estimated gradients $\hat{\mathcal{L}}(P_{n}):=\frac{1}{n}\sum_{i=1}^{n}\hat{\nabla}f\hat{\nabla}f^{T}$. These papers demonstrate, both empirically and theoretically, that this object is closely associated with the performance of simple neural networks, and that its inclusion in a kernel machine allows for estimation adaptive to the global regularity of functions of interest.
(a) Localizations $\mu_{i}$ from Local EGOP Learning (top) and a deep neural network feature embedding (bottom).
(b) Layered EGOP (top) and neural network (bottom) localizations.
Figure 3: Local EGOP Learning and a deep transformer architecture applied to data sampled from an annulus. Labels are generated with no dependence on the radius in the parameterization, with these values overlaid as a heatmap. About the point $x^{*}=(1,0)$, points are displayed with opacity $w\propto\exp(-d(x^{*},x_{i})^{2})$, where $d(\cdot,\cdot)$ is the AGOP Mahalanobis distance (top) and transformer feature embedding (bottom). Both regions are displayed after progressively many iterations/training batches.
However, $\mathcal{L}(\mu)$ may generically be full rank, making procedures such as the aforementioned subject to the curse of dimensionality. The supervised noisy manifold hypothesis is exactly such an example. Generically, $\cup_{x\in\mathcal{M}}\mathcal{T}_{x}=\mathbb{R}^{D}$, the full ambient space, hence globally all subspaces may be informative. Only locally will the EGOP degenerate, necessitating $\mu$ appropriately concentrate about the point of interest $x^{*}$. We study this through the lens of Gaussian localizations, $\mu=N(x^{*},\Sigma)$. The covariance $\Sigma$ and the Mahalanobis distance $M$ have inverted proclivities, as where $M$ seeks to degenerate along the normal, $\mathcal{L}(\mu)$ is most degenerate for $\Sigma$ aligned with the normal. This is no coincidence, as upon selecting $k$ Gaussian, smoothing with weights $\propto k(M^{1/2}(X_{i}-x^{*}))$ corresponds to an average with respect to $\mu=N(x^{*},M^{-1})$.
We formally link these objects in Section 3, developing a rigorous framework to assess the contribution of anisotropy to learning rates for kernel estimators. In sections 4 and 5, we demonstrate how optimizing the corresponding bias-variance trade-off naturally leads to the Local EGOP Algorithm, a recursive procedure reminiscent to the methods proposed in Takeda et al. (2007); Radhakrishnan et al. (2022). Finally, in section 6, we show that this method leads to adaptive regression guarantees for both noisy manifold and generic data. We support the theoretical results with numerical examples illustrating our claims in Section 7, and discuss further directions of research in Section 8.
2 Assumptions and Notation
We refer to the population distribution of our feature vectors $X\in\mathbb{R}^{D}$ as $P$, and its empirical distribution as $P_{n}$. We assume $P$ has bounded $C^{1}$ density with respect to the Lebesgue measure on $\mathbb{R}^{D}$. When invoking the noisy manifold hypothesis, we denote by $\mathcal{M}$ the base manifold of dimension $d$, by $\pi$ the nearest point projection onto the manifold, and by $\mathcal{T}_{x}$ and $\mathcal{N}_{x}$ the tangent and normal spaces, respectively, at point $p:=\pi(x)\in\mathcal{M}$. For $\pi$ to be well-defined, we further assume that $X$ lies within the reach of $\mathcal{M}$ Federer (1959) almost surely. The labels $Y$ are assumed to be of the form $Y_{i}=f(X_{i})+\varepsilon$ for $\varepsilon_{i}$ iid, mean zero, and finite 4th moment. We assume that $f\in C^{4}$, bounded with uniformly bounded derivatives. We reserve $\mu$ to refer to a localization or target distribution of data appropriately concentrated about the point of interest $x^{*}$ with covariance $\Sigma$. We assume a uniform cap for the covariance, $\|\Sigma\|\leq\zeta<\infty$. We denote by $\langle\cdot,\cdot\rangle_{F}$ the Frobenius inner product. This work centers around the EGOP, $\mathcal{L}(\mu):=\int\nabla f(x)\nabla f(x)^{T}\,d\mu$, and its empirical counterpart the AGOP, $\hat{\mathcal{L}}(w):=\sum_{i}w_{i}\hat{\nabla}f\hat{\nabla}f^{T}$, for some estimator $\hat{\nabla}f$ of the gradient. We assume $g=\nabla f(x^{*})\neq 0$. By $k$, we denote a second-order kernel, satisfying typical assumptions in the nonparamtric statistics literature, see for example Tsybakov (2009)[Chapter 1]. In Section 3, we specialize results to the Gaussian kernel $k(\|x\|)\propto\exp(-\|x\|^{2})$, but we discuss the general setting in Appendix B.1.4.
3 Bias Control via EGOP
In this section, we develop a theoretical framework for adaptive kernel machines. Define the empirical kernel estimator for predicting the value of $f$ at $x^{*}$ by
$\hat{P}_{M}(f)$ $:=\frac{1}{\hat{C}_{M}}\frac{1}{n}\sum_{i=1}^{n}k(\|M^{1/2}[X_{i}-x^{*}]\|/\sqrt{2})Y_{i},$ (1)
$\hat{C}_{M}$ $:=\frac{1}{n}\sum_{i=1}^{n}k(\|M^{1/2}[X_{i}-x^{*}]\|/\sqrt{2}),$ (2)
where the $D\times D$ matrix $M\succeq 0$ is a Mahalanobis metric to be learned. Its continuous counterpart is
$P_{M}(f)$ $:=\frac{1}{C_{M}}\int k(\|M^{1/2}[y-x^{*}]\|/\sqrt{2})f(y)\,dP(y),$ (3)
$C_{M}$ $:=\int k(\|M^{1/2}[y-x^{*}]\|/\sqrt{2})\,dP(y).$ (4)
To assess the quality of this predictor, we can measure its bias $\int(P_{M}(f)-f)^{2}\,d\mu$ on a target distribution $\mu$. Our key insight is that, when $M$ and $\mu$ are chosen in correspondence, the bias can be related to a Dirichlet form in terms of the EGOP. To make this explicit, we introduce the EGOP-form
$W(\mu):=\int\nabla f^{T}\Sigma(\mu)\nabla f\,d\mu=\operatorname{tr}(\mathcal{L}(\mu)\Sigma(\mu)).$ (5)
Lemma 1 (Poincaré Inequality).
Let $M_{t}\succeq 0$, and set $\mu_{t}=N(0,M_{t}^{-1})$. Then,
$\int(P_{M_{t}}(f)-f)^{2}\,d\mu_{t}=O(W(\mu_{t})).$
Thus, if the anisotropic metric $M_{t}$ is such that the variation of $f$ around $x^{*}$, given by $\mathcal{L}(\mu_{t})$, and the variance of $x$ itself, given by $\Sigma(\mu_{t})$, are close to low-rank and orthogonal, then the bias of the kernel estimator (3) is controlled by $W(\mu_{t})=\langle\mathcal{L}(\mu_{t}),\Sigma(\mu_{t})\rangle_{F}$, and will be nearly 0. Of critical importance are the first and second moments of the target distribution $\mu$, motivating our Gaussian convention, although Gaussianity is not essential to achieve the desired bound. We note that this general structure, of metrizing by the inverse covariance, appears in disparate literatures, from classical local feature estimators (Schmid and Mohr, 1997), to VAEs (Chadebec and Allassonnière, 2022), and Langevin preconditioning (Cui et al., 2024).
We now seek to control the variance of the finite sample estimator $\hat{P}_{M}$ of $f$ at $x^{*}$.
Lemma 2 (Variance).
Let $M_{t}\succeq 0$ and set $\Sigma_{t}=M_{t}^{-1}$, $\mu_{t}=N(x^{*},\Sigma_{t})$, . Then $\mathbb{E}\left[(\hat{P}_{M_{t}}(f)-P_{M_{t}}(f))^{2}\right]=O\left(1/(n\sqrt{\det\Sigma_{t}})\right).$
Combining these bounds yields our fundamental error rate.
Theorem 1 (Local MSE).
Let $M_{t}\succeq 0$, and set $\Sigma_{t}=M_{t}^{-1}$, $\mu_{t}=N(x^{*},\Sigma_{t})$. Then,
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t}\right]=O\left(\frac{1}{n\sqrt{\det\Sigma_{t}}}+W(\mu_{t})\right).$
In particular, if $W(\mu_{t})=O(t)$ and $\det\Sigma_{t}=O(t^{q})$ then the asymptotic learning rate is
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t}\right]=O\left(n^{-\frac{1}{1+q/2}}\right).$
Thus, if we design a sequence $\{t_{i}\}_{i\in\mathbb{N}}$, with $t_{i}\to 0$, such that $M_{t_{i}}$ and $\mu_{t_{i}}$ are updated recursively based on their values up to $t_{i-1}$ and $\frac{1}{n\sqrt{\det\Sigma_{t_{i_{n}}}}}+W(\mu_{t_{i_{n}}})\to 0$, the kernel estimator $\hat{P}_{M_{t_{i_{n}}}}$ is consistent and its rate is controlled by the convergence rate of $W$ and $M$.
4 The Local EGOP Algorithm
In this section, we introduce the Local EGOP Learning algorithm, detailed in Algorithm 1. The theoretical motivation for this iterative scheme is presented in Section 5. The algorithm iteratively metrizes a kernel by an AGOP, leveraging this operator to assign weights to points neighboring $x^{*}$.
Input: Data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, target $x^{*}$, subsample size $m$, iterations $T$, bandwidths $\{t_{i}\}$, $\beta>0$, scale $\alpha>0$.
Output: Best estimate $\widehat{f}(x^{*})$.
Hyperparameters: $M_{0}\leftarrow I/\alpha$.
1
[…]
3
4for $i=0$ to T1T-1 do
 $\triangleright$ Compute global weights at $x^{*}$
5 $w_{j}\propto\exp\!\bigl(-(X_{j}-x^{*})^{\top}M_{i}(X_{j}-x^{*})\bigr),\ j=1,\dots,n$;
6   Normalize $w$, subsample $S\subset\{1,\dots,n\},|S|=m$ with probs $w$;
7
 $\triangleright$ Leave-One-Out Local regressions at each $j\in S$
8 for $j\in S$, do
9      Fit weighted local linear regression centered at $X_{j}$ (weights from metric $M_{i}$), data point $(X_{j},Y_{j})$ excluded;
10      Obtain $\widehat{\nabla}f[j]$ and $\widehat{f}[j]$;
11
12
 $\triangleright$ Aggregate statistics
13 $L_{i}\leftarrow\sum_{j\in S}w_{j}\,\widehat{\nabla}f[j]\widehat{\nabla}f[j]^{\top}$;
14 $\widehat{\mathrm{MSE}}_{i}\leftarrow\sum_{j\in S}w_{j}\,(Y_{j}-\widehat{f}[j])^{2}$;
15
16 if $\widehat{\mathrm{MSE}}_{i}<\text{bestMSE}$ then
17      Estimate $\widehat{f}(x^{*})$ via local regression centered at $x^{*}$;
18    $\text{bestMSE}\leftarrow\widehat{\mathrm{MSE}}_{i}$,  $\widehat{f}_{\text{best}}(x^{*})\leftarrow\widehat{f}(x^{*})$;
19
20
 $\triangleright$ Metric update: set $\beta=1$ if $i=0$, $\beta<1$ otherwise
21 $M_{i+1}\leftarrow\bigl(\beta L_{i}+(1-\beta)L_{i-1}\bigr)$;
22 $M_{i+1}\leftarrow M_{i+1}/[t_{i+1}\operatorname{tr}(M_{i+1})]$;
23
24
return $\widehat{f}_{\text{best}}(x)$
Algorithm 1 Local EGOP Learning
The time complexity of the method is $O(T((m+2)nD^{2}+(m+1)D^{3}))$, with memory overhead $O(nD^{2}+m^{2}D^{2}+mD^{3})$.
Incorporated into this algorithm is a one-step lookback (step 13) facilitated through $\beta\in(0,1)$. This is essential to improve stability, as discussed in Example 2. In practice, one might choose hyper-parameters $\beta,T$, as well as the schedule $\{t_{i}\}$ via cross-validation. Additionally, for computational reasons, we normalize $M_{i}$ by its trace (step 14) before applying the scaling factor. This has little impact asymptotically, as the trace converges to the leading gradient eigenvalue of the EGOP, making this normalization effectively equivalent to normalization by a constant scalar. However, it serves a practical benefit by stabilizing initial iterates, setting the scale of the metric to the predefined rate $1/t_{i}$. Note that while the optimal number of iterations is dimension dependent, as argued in Theorems 1 and 3, our algorithm incorporates a Leave-One-Out style validation loop (step 10) for adaptive stopping.
5 Recursive Kernel Learning as Variance Minimization
We will now show how recursive kernel learning, combined with our basic kernel bounds, constitutes a greedy variance reduction strategy. Intuitively, the goal is to induce anisotropy in the kernel, stretching or steering it to include additional low-bias points that would otherwise be discarded with a RBF. The inclusion of these additional high-quality points is exactly a reduction of variance, as including more data results in tighter concentration to the mean for a local averaging estimator.
We now make precise how one can arrive at this by leveraging Theorem 1. Ideally, as spelled out in this result, we would like to select $\mu_{t}$ such that $W(\mu_{t})\to 0$, and that $\det\Sigma(\mu_{t})$ is as large as possible. In other words, if we specify the rate $W(\mu_{t})=\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]=t\to 0$, it would be optimal to select the largest possible region that achieves this bias, meaning we would like to select
$\mu_{t}=\operatorname{argmax}_{\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq t}\log\det\Sigma(\mu_{t}).$
This, however, cannot be solved, as we lack precise knowledge of the function of interest $f$. Moreover, $\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq t$ presents a challenging feasible set, as it is posed over a space of probability distributions, thus making it non-Euclidean. Our Gaussian ansatz reduces the problem to optimizing a finite parameterization, however this still presents a non-convex optimization.
This leads us to the following discretized relaxation of the problem. Let $t_{i}\to 0$ and set $\mu_{0}=N(x^{*},\alpha I)$, an isotropic initialization. If we compute the initial $\hat{\mathcal{L}}(\mu_{0})$, we may hope that, as long as $\mu_{1}$ represents a sufficiently comparable distribution, $\hat{\mathcal{L}}(\mu_{0})\approx\hat{\mathcal{L}}(\mu_{1})$. Thus, we can compute
$\Sigma_{1}=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{0})\Sigma]\leq t_{1}}\log\det\Sigma=t_{1}\hat{\mathcal{L}}(\mu_{0})^{-1}/D.$
This yields $\mu_{1}=N(x^{*},\Sigma_{1})$, and the corresponding kernel estimator is metrized by $\Sigma_{1}^{-1}\propto\hat{\mathcal{L}}(\mu_{0})$. Repeating this, we have a recursive kernel learning algorithm, as we summarize below.
Proposition 1 (Alternating Minimization).
Let $t_{i}\to 0$, fix $\Sigma_{0}$, $\mu_{i}=N(x^{*},\Sigma_{i})$, $\hat{\mathcal{L}}_{i}:=\hat{\mathcal{L}}(\mu_{i}):=\int\hat{\nabla}f\hat{\nabla}f^{T}\,d\mu_{i}$, and construct $\hat{\nabla}f=\hat{P}_{\Sigma_{i}^{-1/2}}(f)$. Define the recursive kernel learning algorithm $\Sigma_{i+1}=t_{i+1}\hat{\mathcal{L}}_{i}^{-1}/D$. Then, equivalently,
$\Sigma_{i+1}$ $=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]\leq t_{i+1},\Sigma\succeq 0}\log\det\Sigma$ (6)
$=\operatorname{argmin}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]=t_{i+1},\Sigma\succeq 0}\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]+\frac{1}{n\sqrt{\det\Sigma}}.$
We note that in Algorithm 1, we use local linear regression to better fit into the EGOP estimation loop. We expect that a suitable extension of our theory is possible for higher-order polynomials, allowing for more complex local regressions in Algorithm 1. See Section 8 for further details.
From this framework we have developed, it is clear that there are many possible schemes to optimize the bias–variance trade-off, and we discuss this in more detail in Section 8. We choose to focus on this alternating optimization scheme because it strongly resembles algorithms present in the existing literature, and is simple to implement.
6 Guarantees
In this section, we demonstrate concrete theoretical guarantees for estimation via Local EGOP Learning. In our theory, we do not consider the problem of EGOP estimation, and instead consider an iterative scheme in which we can query an oracle for the EGOP associated to a particular choice of covariance. For our experiments, we compute empirical AGOPs fit via the proposed iterative estimation procedure, leading to sharp estimation quality. We leave it to future work to close this theoretical gap, although we discuss it in greater depth in Appendix C. See Appendix B for proofs of all following arguments.
6.1 Taylor Expansion
Up to this point, our prevailing perspective is that the target distributions $\mu$ matter up to their first and second moments, passing the optimization problem from the space of measures to the PSD cone, and informing a Gaussian relaxation. We close the loop to EGOP estimation by additionally Taylor expanding $\mathcal{L}(\mu)$ up to second order, reducing the proposed iteration scheme entirely to a recurrence in $\Sigma$.
Lemma 3 (Generic Taylor Expansion).
Let $\mu=N(x^{*},\Sigma)$, $g=\nabla f(x^{*})$, and $H=\nabla^{2}f(x^{*})$. Then there exist $T:\mathbb{R}^{D\times D}\to\mathbb{R}^{D}$, $R:\mathbb{R}^{D\times D}\to\mathbb{R}^{D\times D}$, $C_{T}\geq 0,$ and $C_{R}\geq 0$ such that for any $A\succeq 0$, $\|T(A)\|\leq C_{T}\|A\|,\ \|R(A)\|\leq C_{R}\|A\|^{2}$, and
$\mathcal{L}(\mu)=gg^{T}+H\Sigma H+gT(\Sigma)^{T}+T(\Sigma)g^{T}+R(\Sigma).$
Define $T_{g}(\Sigma):=gT(\Sigma)^{T}+T(\Sigma)g^{T}$. Our goal is to analyze
$\Sigma_{i+1}$ $=t_{i+1}(\beta\mathcal{L}(\mu_{i})+(1-\beta)\mathcal{L}(\mu_{i-1}))^{-1}$ (7)
$=t_{i+1}(gg^{T}+\beta[H\Sigma_{i}H+T_{g}(\Sigma_{i})+R(\Sigma_{i})],$
$\quad\quad+(1-\beta)[H\Sigma_{i-1}H+T_{g}(\Sigma_{i-1})+R(\Sigma_{i-1})])^{-1}.$
We initialize this recurrence with the isotropic choice $\Sigma_{0}=\alpha I$, $\Sigma_{1}=t_{1}\mathcal{L}(\mu_{0})^{-1}$, $\Sigma_{2}=t_{1}\mathcal{L}(\mu_{1})^{-1}$, then apply momentum following this step.
6.2 Full Rank Hessian
We begin our analysis with the generic setting, where $H$ is full-rank. In this setting, the Taylor expansion of $\mathcal{L}(\mu)$ is dominated by a rank 1 matrix $gg^{T}$ and a linear term $H\Sigma H$, which combine to yield a full rank matrix. We show that Local EGOP Learning results in a marked efficiency gain over generic isotropic kernel methods. Indeed, we argue that $g^{T}\Sigma_{i}g=\Theta(t_{i})$, while for $v\perp g,$ $v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$, thus achieving a second-order anisotropy. The noisy manifold setting implies a rank deficiency of $H$, and we will show in the next section how this leads to further efficiency gains.
Example 1 (Scalar Recurrence).
Fix a schedule for $t_{i}\to 0$ that is no faster than geometric, so that $t_{i+1}/t_{i}$ is larger than a fixed threshold. We investigate a considerable reduction of the problem, the one dimensional recurrences
$a_{i+1}$ $=\frac{t_{i+1}}{\beta a_{i}+(1-\beta)a_{i-1}},$ (8)
$b_{i+1}$ $=\frac{t_{i+1}}{c+\beta b_{i}+(1-\beta)b_{i-1}}.$ (9)
It is easy to see that $b_{i}=O(t_{i})$, as if $b_{i}$ decays at all, then we have $b_{i}=t_{i}/c+o(t_{i})$. One can also readily observe, at least informally, the rate of decay of $a_{i}$. Assume that $a_{i}=\Theta(t_{i}^{\zeta})$. By the recurrence relation, we then have $a_{i}=\Theta(t_{i}^{1-\zeta})$, hence these coincide at the second order anisotropy $\alpha=1/2$. As our matrix recurrence is primarily given by a homogeneous part $(H\Sigma H)$ plus a rank deficient constant part $(gg^{T})$, we argue in Appendix B that the $b_{i}$ correspond well to the $g\times g$ block of $\Sigma_{i}$, and $a_{i}$ to $g^{\perp}\times g^{\perp}$.
Example 2 (Momentum).
We now take a moment to highlight the necessity of the momentum term $\beta>0$. Consider, again, the sequence $a_{i}$ from Example 1, and suppose we set $\beta=1$. For convenience, let us assume $\sqrt{t_{i+1}}=\sqrt{t_{i}}r$, an exact geometric schedule. We reparameterize, writing
$a_{i+1}=\frac{t_{i+1}}{a_{i}}\Longleftrightarrow a^{\prime}_{i+1}=\frac{1}{a^{\prime}_{i}}$
for $a_{i}^{\prime}:=\frac{a_{i}}{\sqrt{rt_{i}}}$. Clearly, $a_{i}^{\prime}$ has no limit, as it simply oscillates wildly about $1$. The remainder terms, as well as the estimation error, make the $\sqrt{t_{i}}$ rate become unstable, leading to potential catastrophic failure. We demonstrate the resulting localizations with and without momentum in Figure 4.
Figure 4: Comparison of the eigenvalue decay of $\Sigma_{i}$ for Local EGOP Learning with and without momentum. On the left, we set $\beta=0.7$, on the right $\beta=1$ (no momentum). Without momentum, the second order eigenvalues are prone to wild oscillations.
This leads to the following result.
Theorem 2 (Second Order Anisotropy).
Let $H$ be invertible and $\Sigma_{i}$ satisfy the recurrence in equation 7, $\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, and for $v\perp g$, $v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$. It follows that, selecting $i_{n}=\frac{4\log n}{(D+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{\Sigma_{i_{n}}}(f)-f)^{2}\,d\mu_{i_{n}}\right]=O\left(n^{-\frac{4}{D+5}}\right).$
Compared to a vanilla 0-order local polynomial estimator, this method nearly squares the asymptotic error, effectively cutting the effective dimension in half for $D$ large. In our theory, we assume the initial localization has small covariance $\alpha I$. In the resulting application of Lemma 3, this results in an explicit recurrence up to a higher-order remainder. While a useful theoretical construct, this is not necessary empirically, and we leave it to future research for a complete analysis of the iteration stability.
6.3 Noisy Manifold
In the noisy manifold setting we strengthen Lemma 4.
Lemma 4 (Noisy Manifold Taylor Expansion).
Let $g=\nabla f(0)$, $H=\nabla^{2}f(0)$, $p=\pi(0)$, $\mathcal{T}:=\mathcal{T}_{p},\mathcal{N}:=\mathcal{N}_{p}$, and $T,R$ be as in Lemma 4. There exist non-negative constants $C_{\mathcal{N}}$ and $C_{\mathcal{T}\mathcal{N}}$ such that, for all $A\succeq 0$,
$\|\pi_{\mathcal{N}}R(A)\pi_{\mathcal{N}}\|$ $\leq C_{\mathcal{N}}\|\pi_{\mathcal{T}}A\pi_{\mathcal{T}}\|^{2},$ (10)
$\|\pi_{\mathcal{T}}R(A)\pi_{\mathcal{N}}\|$ $\leq C_{\mathcal{T}\mathcal{N}}\|\pi_{\mathcal{T}}A\pi_{\mathcal{T}}\|.$ (11)
Additionally, $\operatorname{rank}(H)\leq 2d.$
We will leverage this weak orthogonal dependence to show that the distribution $\mu$ elongates along the space $\mathcal{N}_{x^{*}}$ orthogonal to the manifold. However, while the function is constant on this fiber, its gradient is not. Indeed, we are only guaranteed that the gradient remains tangent. Due to the effects of curvature, $\nabla f(p+v)$ may shrink or grow along different coordinate directions. We characterize this precisely as follows.
Lemma 5 (Gradient Geometry).
Let $p\in\mathcal{M}$, $\eta\in\mathcal{N}_{p}$, $\|\eta\|<\tau$, $\tau$ the reach of $\mathcal{M}$. Then,
$\nabla f(p+\eta)=(I-S_{p}(\eta))^{-1}\nabla f(p).$
In particular, for any such $\eta$, there exists a subspace $\operatorname{Null}=\operatorname{Null}_{\eta}$ of dimension at least D2dD-2d such that for any $v,z\in\operatorname{Null}$, $\nabla f(p+\eta+v)=\nabla f(p+\eta)$, $\nabla^{2}f(p+\eta+v)z=0$.
These geometric implications allow us to further adapt Lemma 4 to our geometry.
Corollary 1 (Shifted Taylor).
Let $x^{*}=p+\eta$, $v\in\operatorname{Null}_{\eta}$. Define $\mu_{v}$ as the measure $\mu=N(x^{*},\Sigma)$ conditioned on $\pi_{\operatorname{Null}}[X-x^{*}]=v$, i.e. the draws where the $\operatorname{Null}$ component is fixed at $v$. Let $\Sigma_{v}$ be the covariance of this distribution. Then,
$\mathcal{L}(\mu_{v})=gg^{T}+H_{v}\Sigma_{v}H_{v}+gT_{v}(\Sigma_{v})^{T}+T_{v}(\Sigma_{v})g^{T}+R_{v}(\Sigma_{v})$
where $\operatorname{Null}\subseteq\operatorname{nullspace}(H_{v})$, $\sup_{v\in\operatorname{Null}}\|R_{v}(A)\|\leq C_{\operatorname{Null}}\|A\|^{2},\ T_{v}(A)\leq C_{\operatorname{Null}}\|A\|$ for a uniform constant $C_{\operatorname{Null}}$, and $H_{v},T_{v},R_{v}$ are continuous in $v$, Lipschitz in $\|A\|$.
This allows us to perform an iterated integration, freezing the Taylor expansion along this shifted base point while preserving the fundamental first order behavior. To simplify our argument we assume that the covariance cap $\zeta$ is small, allowing a simple reduction of the dynamics to the non-degenerate case. This greatly streamlines the analysis, although we do not believe it is fundamental to the convergence.
Theorem 3 (Intrinsic Learning Rate).
Assume that $D\geq 2d$ and $\operatorname{rank}(H)=2d$. Let $\Sigma_{i}$ satisfy the recurrence in equation 7, $\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, for $v\in\operatorname{Null}$, $v^{T}\Sigma_{i}v=\Theta(1)$, and for $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\theta(\sqrt{t_{i}})$. It follows that, selecting $i_{n}=\frac{4\log n}{(2d+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{\Sigma_{i_{n}}}(f)-f)^{2}\,d\mu_{i_{n}}\right]=O\left(n^{-\frac{4}{2d+5}}\right).$
Rather than fully denoising the manifold, our analysis reveals that the manifold tangent, as well as the normal vectors that do not fall in $\operatorname{Null}$, are subject to the same second-order anisotropy as observed in the full rank setting. These compose precisely those directions in which the gradient varies at first order, leaving only the inactive subspace where the linearization is preserved. Practically, the resulting rate is comparable to that of a standard estimator applied to a completely denoised manifold, as this method yields an effective dimension of $d+1/2$.
Example 3 (Spheres).
Built into Theorem 3 are assumptions on the ambient dimension and derivatives of $f$. The sphere $S^{D-1}$ is a typical toy example, however $D\geq 2(D-1)$ fails for $D\geq 3$. In these settings, the hessian has the form
$H=\begin{bmatrix}T&b\\ b^{T}&0\end{bmatrix}$
in the $\mathcal{T}\times\mathcal{N}$ basis. Thus the noisy manifold hypothesis only guarantees that a single entry of the Hessian is 0. Hence, $\pi_{g^{\perp}}H\pi_{g^{\perp}}$ is generically full rank, driving second order anisotropy along the tangent and orthogonal, as demonstrated in Figure 5.
Example 4 (Continuous Single-Index).
When the underlying manifold is 1-dimensional, the tangent space coincides with the span of the gradient, hence $\pi_{g^{\perp}}H\pi_{g^{\perp}}=0$, removing the second order anisotropy. Thus no normal directions contract asymptotically, leading to an effective dimension of 1 for the learning rate. This corroborates the result of Wu and Maggioni (2024).
Figure 5: Eigenvalue decay of $\Sigma_{i}$ with Local EGOP Learning applied to data satisfying the noisy manifold hypothesis about $S^{2}$. The orthogonal exhibits light decay, approaching second order anisotropy asymptotically.
Example 5 (Multi-Index Learning).
When the underlying manifold $\mathcal{M}$ is flat, the supervised noisy manifold hypothesis reduces to a multi-index setting. The key distinction is that displacement away from the manifold not only preserves the function value, but also all higher order function derivatives. In particular, the Hessian has rank no more than $d$, the active index dimension, and the Hessian column space is shared across the distribution.
Thus, when performing expansions such as Corollary 1, the orthogonal has no contribution to the EGOP, leading to implicit dimension reduction. In particular, this yields a learning rate of $(d+1)/2$.
7 Experiments
We provide the following simulations to empirically justify our theoretical results. First, expanding on the noisy manifold setting, we show how the regression error rates of Local EGOP Learning remain invariant to the injection of high-dimensional noise. We demonstrate that two-layer neural networks are not able to efficiently learn in the noisy manifold setting, with a sharp decrease in performance compared to Local EGOP Learning. Next, we compare the performance of our method to that of a deep neural network on toy data. In particular, we show how the feature embeddings produced by transformers trained on a simple example are qualitatively similar to the localizations generated by our procedure. Finally, we apply this procedure to estimate backbone angles in Molecular Dynamics (MD) data, where we leverage the noisy manifold structured data to improve prediction quality.
In our simulations, we consider helical data, parameterized by a curve $\theta(t)=(\sin(t+w_{1}),\cos(t+w_{1}),\sin(t+w_{2}),\cos(t+w_{2}),\dots,g(t))$, where $g(t)=t$ is a linear term included if $D$ is odd dimensional, and the $w_{i}$ are constant offsets taken as a mesh from 0 to $2\pi$. We rescale this data by a constant $\tau$, then contaminate it with uniform orthogonal noise of radius $r$. In our simulations, we set $\tau=0.8,\,r=0.5$, and sample $t$ from 0 to $2\pi$. See Figure 7(a) for a visualization. For the outcomes $y$, we generate a third-degree polynomial with coefficients uniformly sampled from (3,3)(-3,3), then evaluate it at the projection point onto $\theta(t)$.
We relegate details for the evaluation and fitting of each method to Appendix A.
Figure 6: (Left) Comparison of Local EGOP Learning to the performance of a two-layer neural network architectures trained on helical data in ambient dimension $D=5$. (Right) Local EGOP Learning trained on helical data in various ambient dimensions $D$.
7.1 Learning Rate
We generate helical data in a variety of dimensions, testing the Local EGOP Learning algorithm. As seen in Figure 6, the asymptotic learning rate has little dimensional dependence. However, error scaling depends strongly on whether $D$ is even or odd, and this is because the underlying manifold changes in these settings, although its intrinsic dimension does not.
7.2 Two-layer Neural Network
We assess a variety of two-layer neural network architectures, with implementation and training explained in detail in Appendix A.
Helical data provides a rich experimental setting to compare Local EGOP Learning to two-layer neural networks. It is not only continuous single-index, it is multi-index 2 for $D$ even, and 3 for $D$ odd. Further, by construction it is well-approximated by a low intrinsic dimensional manifold, a typical setting for accelerated learning rates (Kiani et al. (2024); Liu et al. (2021), etc.).
Unsurprisingly, as demonstrated in Figure 6, the two-layer neural networks readily adapt to the multi-index data structure. For $D=5$, the overparameterized ReLU network achieves comparable error to a multi-index architecture pre-specified to learn a 3-dimensional subspace. However, Local EGOP Learning vastly reduced the error for moderate to large sample sizes. This suggests that no architecture was able to fully adapt to the intrinsic continuous single-index structure.
7.3 Feature Learning
In this section, we compare the local feature learning capabilities of a deep transformer-based neural network (Gorishniy et al., 2023) to Local EGOP Learning. We consider data generated from a 1-sphere under the supervised noisy manifold hypothesis. On this simple dataset, both algorithms result in nearly complete denoising of the data, as shown in Figures 3(a) and 3(b). In Appendix A, we provide additional unsupervised embeddings for comparison.
7.4 Predicting the backbone angles in Molecular Dynamics (MD) data
Our final example comes from the analysis of molecular geometries. The raw data consist of $X,Y,Z$ coordinates for each of the $N_{a}$ atoms of a molecule, which, due to interatomic interactions, lie near a low-dimensional manifold (Das et al., 2006). While the governing equations of the simulated dynamics are unknown, for small organic molecules, certain backbone angles (Das et al., 2006) have been observed to vary along the aforementioned low-dimensional manifold. Specifically, for the malonaldehyde molecule, the two backbone angles denoted $\tau_{1,2}$ are shown in Figure 7(b). We used a subsample of molecular configurations of size $n=10^{4}$ from the MD simulation data of Chmiela et al. (2017) as input data. The configuration data, pre-processed as in Koelle et al. (2022), consists of $D=50$ dimensional vectors and lie near a 2-dimensional surface with a torus topology (see Figure 7(b)). On a hold-out set of 500 test points, Local EGOP Learning yields an MSE of 0.00043, compared to 0.012 for Gaussian kernel Nadaraya-Watson with cross-validated bandwidth selection.
8 Discussion and Future Work
The primary contribution of this work is to establish a rigorous basis for the study of recursive kernel machines that localize about points of interest. We leverage this framework to verify that Local EGOP Learning achieves desirable learning rates in model settings. Of particular importance is our introduction of continuous index learning, formalized by the supervised noisy manifold hypothesis, under which anisotropy about the orthogonal is achieved under appropriate non-degeneracy assumptions. Our algorithm is most similar to the heuristic Takeda et al. (2007), hence we believe that the present framework can be extended to analyze this algorithm, too. Local EGOP takes its name from Radhakrishnan et al. (2022), with the key distinction being that the AGOP is averaged over the localized region $\mu_{i}$ rather than the entire data distribution $P_{n}$.
We demonstrate the efficacy of this algorithm, qualitatively when compared to a transformer architecture, and quantitatively on high dimensional continuous single-index data. The theoretical framework developed in this work opens many avenues for future study. In addition to the open problems this investigation has raised, our approach is conducive to a variety of strategies that may further improve the stated results. We mention several of the most relevant directions.
For one, we can consider optimizing the objective
$\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]+\frac{1}{n\sqrt{\det\Sigma(\mu_{t})}}$
in a variety of ways. Working with a second order relaxation, one may seek to optimize the covariance gradually, such as by a gradient-based method. Alternatively, motivated by Proposition 1, one can show that this ratio is minimized by a distribution $\mu_{t}$ such that $\Sigma(\mu_{t})=t\mathcal{L}(\mu_{t})^{-1}/D$. Differentiating, this yields a differential equation
$\partial_{t}\mathcal{L}(\mu_{t})\Sigma(\mu_{t})+\mathcal{L}(\mu_{t})\partial_{t}\Sigma(\mu_{t})=I/D$
which may be amenable to analysis, particularly in the Gaussian setting. However, it may also be fruitful to drop both this assumption and the covariance relaxation. One promising direction is to view $W(\mu_{t})$ and $\Sigma(\mu_{t})$ as functionals, and to then optimize the ratio using the Wasserstein gradient flow (Ambrosio et al., 2006) or other methods that operate at the distribution level. In particular, this would allow for non-ellipsoidal localizations, potentially allowing for improved learning rates in settings where the level-set is curved. A fundamental hurdle for any of these approaches is that, generally, the EGOP-form can only upper bound localization-induced bias, and the resulting bound is not always sharp.
Many basic questions have been left for future research. As discussed in Appendix C, in the context of Local EGOP Learning, the problem of EGOP estimation is left open. Algorithm 1 is slow, requiring many local linear regressions at each iteration. We presented this simple algorithm for clarity, and many practical improvements can be made to improve its performance, particularly when multiple data points are queried for prediction. Furthermore, simple modifications to the algorithm would yield gains in statistical efficiency. Higher-order local polynomial regression can be performed for function estimation. This would improve theoretical guarantees by accounting for more general smoothness patterns in the function of interest, but it may also inform more complex localization procedures. Gaussian localizations are quadratic in nature, and so a more generic formulation may yield increased adaptivity.
Importantly, Local EGOP Learning is based firmly on a MSE formulation of regression, restricting its utility to $\ell_{2}$ loss. It is unlikely that this method can be applied directly to problems such as classification or image generation. This leaves open how an appropriate adaptation should be formulated in these settings. Additionally, a formal analysis comparing procedures such as Local EGOP Learning to modern neural network architectures is merited. In particular, while we do not expect an exact replication of behavior, the Local EGOP ansatz, which asserts that learned features ought to spread locally inversely proportionate to the EGOP, may inform further study of these learning algorithms, or lead to improvements in the design of model architectures.
Acknowledgements
The authors would like to thank Misha Belkin for the initial suggestion of comparing Local EGOP Learning to the performance of two-layer neural networks, and Brennan Dury for suggestions and experiments related to earlier conceptions of this research.
Appendix A Experiment Details
The simulations were run on a machine of Intel® Xeon® processors with 48 CPU cores, and 50GB of RAM.
(a) 3-dimensional Helical data.
(b) Coordinates of Malonaldehyde embedded into 2 dimensions via Diffusion Maps [Coifman and Lafon, 2006] and colored by the backbone angles $\tau_{1}$.
A.1 General Set-up
In our synthetic experiments, to generate labeled data we randomly generated a homogeneous cubic polynomial
$f(x)=\sum_{|\alpha|\leq 3}c_{\alpha}x^{\alpha},$
where $c_{\alpha}\sim\mathrm{Uniform}[-3,3]$. The resulting function $f$ is then evaluated at the unperturbed datapoints before the features $X$ were contaminated with noise orthogonal to the underlying manifold $\mathcal{M}$. The coefficients $c_{\alpha}$ are randomly sampled across each replicate of the experiments. We generically apply the below hyperparameters to the Local EGOP Learning algorithm in each example.
  • number of EGOP iterations $T_{\text{test}}=150$,
  • Monte Carlo sample size $m=300$,
  • Bandwidth decay $t_{i}=(1+i)^{-1.2}$,
  • Initialization scale $\alpha=0.2$
  • Exclusion radius $\varepsilon=1.6$
A.2 Two-layer Neural Network
We train fully connected NNs with both ReLU and GELU activations. All first layers use Kaiming initialization, and output layers are initialized at variance $1/\sqrt{m}$ where $m$ is hidden width. All networks are trained with Adam.
For moderate-scale neural models, we conduct a grid search over hyperparameters. Widths are selected from $128,256,512,1024,2048,4096,8192$, learning rates are selected from $10^{-3},3\times 10^{-3}$, weight decay from $0,10^{-5},10^{-4}$, and batch size from $16,32,64$. The default Adam momentum parameters are used. Early stopping is applied with patience $200$ and improvement tolerance $10^{-4}$ monitored on a validation split. In each experiment, only the optimal test error of these methods is aggregated into the “Small Tuned NN” plot.
We train overparameterized two-layer models with width $100,000$, Adam with learning rate $0.003$, batch size $64$, and fixed weight decay $10^{-4}$, applying early stopping with patience $300$ and improvement tolerance $10^{-4}$.
We also consider SIL models $f(x)=g(\beta^{\top}x)$, $g$ a two-layer NN, with width 256 and GELU activations, $\beta$ is normalized at each iteration.
A.3 Feature Learning
To qualitatively compare the local weighting behavior of EGOP with that of a learned neural embedding, we constructed a unified visualization pipeline producing animated opacity maps over a shared dataset. Both models were applied to the same two-dimensional manifold samples, with visualization frames showing how point-wise influence weights evolve during training or iteration.
Setup.
We generated n=50,000n=50{,}000 samples on a spherical cap in $\mathbb{R}^{2}$, restricted to an angular width of $25^{\circ}$. Each point $x_{i}$ was scaled radially by a random factor $r_{i}\sim\mathrm{Uniform}[r,R]$, with $r=0.6$ and $R=1.8$, and colored by the output of a random cubic polynomial $f(x)$. Additive Gaussian noise with standard deviation $0.05$ was applied to the responses. The cap’s pole $x_{0}=(1,0)^{\top}$ was used as a fixed reference point for visualization of weight evolution.
We trained an FTTransformer regression model [Gorishniy et al., 2021] on this dataset using the rtdl and zero frameworks. The network was trained for 40 epochs with a batch size of 1024 and learning rate $2\times 10^{-3}$. After every 50 training steps, we computed an embedding-based influence profile analogous to EGOP weights:
$w_{i}=\frac{\exp(-\|z_{i}-z_{0}\|^{2}/\tau)}{\sum_{j}\exp(-\|z_{j}-z_{0}\|^{2}/\tau)},$
where $z_{i}$ denotes the learned feature representation of $x_{i}$, $z_{0}$ is the embedding of the reference point, and $\tau=8.0$ is a temperature parameter controlling decay. These weights were rendered with the same opacity scheme as EGOP to enable direct visual comparison.
Appendix B Proofs
In our theory, we focus on the kernel $k=\exp(-\|x\|^{2})$, as this allows for a direct correspondence to the risk distribution $\mu$ and its integration operators. We discuss general kernels in Appendix B.1.4.
B.1 Bias Control via EGOP
B.1.1 Bias Bound
To derive the upper bound from Lemma 1, we argue in stages. We first decompose the kernel smoother into two parts. The first is given by a pure Gaussian convolution and a density dependent covariance relating $f$ and $p$. Both terms can readily be controlled via the EGOP-form, leading to our bound.
Define $\mu_{M}:=N(x^{*},M^{-1}),$ $\mathbb{E}_{M}[\cdot]:=\mathbb{E}_{\mu_{M}}[\cdot]$, $\operatorname{Cov}_{M}$, $\operatorname{Var}_{M}$ similarly.
Lemma 6 (Basic Decomposition).
$P_{M}f=\mathbb{E}_{M}f+\frac{\operatorname{Cov}_{M}[f,p]}{\mathbb{E}_{M}p}=:G_{M}f+A_{M}f$
Proof.
We can simply compute
$P_{M}f$ $=\frac{\mathbb{E}_{M}[fp]}{\mathbb{E}_{M}[p]}=\frac{\mathbb{E}_{M}[f]\mathbb{E}_{M}[p]}{\mathbb{E}_{M}[p]}+\frac{\operatorname{Cov}_{M}[f,p]}{\mathbb{E}_{M}[p]}$
$=\mathbb{E}_{M}[f]+\frac{\operatorname{Cov}_{M}[f,p]}{\mathbb{E}_{M}[p]}=:\mathbb{E}_{M}[f]+A_{M}f.$
Lemma 7 (Gaussian Poincaré).
$\int(\mathbb{E}_{M}[f]-f)^{2}\,d\mu_{M}\leq W(\mu_{M})$
Proof.
This is simply the Gaussian Poincaré inequality, see for instance Bakry et al. [2013][Proposition 4.1.1]. ∎
Lemma 8 (Poincaré $\times$ $p$ variance).
$(A_{M}f)^{2}\leq W(\mu_{M})\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}$
Proof.
$A_{M}f^{2}$ $=\frac{\operatorname{Cov}_{M}[f,p]^{2}}{\mathbb{E}_{M}[p]^{2}}\leq\frac{\operatorname{Var}_{M}(f)\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}\leq W(\mu_{M})\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}},$
applying the Gaussian Poincaré inequality Bakry et al. [2013][Proposition 4.1.1] for the last line. ∎
Proof of Lemma 1.
We can apply Lemmas 7 and 8 immediately to yield
$\int(P_{M}-f)^{2}\,d\mu_{M}=\int(\mathbb{E}_{M}[f]-f+A_{M}f)^{2}\,d\mu_{M}=\int(\mathbb{E}_{M}[f]-f)^{2}\,d\mu_{M}+A_{M}f^{2}\leq\left(1+\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}\right)W(\mu_{M}).$
Note that the density dependent factor $\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}$ is bounded by the continuity of $p$ and the capped covariance assumptions.
B.1.2 Variance Bound
We operate by a typical Nadaraya-Watson style argument. Our first steps are to show that the denominator of our estimator is very concentrated, allowing us to treat it as a constant up to a negligible residual. Let $\gamma_{M}\propto\det(M)^{1/2}$ by the normalization factor of $\mu_{M}$. Define
$\hat{C}^{\prime}_{M}:=\gamma_{M}\hat{C}_{M}:=\gamma_{M}\frac{1}{n}\sum_{i=1}^{n}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*})).$
Lemma 9 (Denominator Concentration).
$\mathbb{P}(|\hat{C}^{\prime}_{M}-\mathbb{E}_{M}(p)|\leq\mathbb{E}_{M}(p)/2)\leq 2\exp\left(-\frac{n\mathbb{E}_{M}(p)}{10\gamma_{M}}\right)$
Proof.
We prove this by applying Bernstein’s inequality. First, observe $\mathbb{E}[\hat{C}_{M}]=G_{M}(p)$. Next, $\exp(-v^{T}Mv)\in(0,1]$ for any $v$, hence $\hat{C}^{\prime}_{M}=\frac{\gamma_{M}}{n}\sum_{i=1}^{n}\exp(-(X_{i}-x)^{T}M(X_{i}-x))$ is the sample mean of iid bounded random variables in $(0,\gamma_{M}]$. Further, we can compute
$\operatorname{Var}[\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x))]\leq\gamma_{M}\mathbb{E}[\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x))^{2}]$
$\quad\quad\leq\gamma_{M}\mathbb{E}[\exp(-(X_{i}-x)^{T}M(X_{i}-x))]=\gamma_{M}G_{M}(p).$
Hence,
$\mathbb{P}(|\hat{C}^{\prime}_{M}-G_{M}(p)|$ $\leq-G_{M}(p)/2)\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\operatorname{Var}(\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x)))+\frac{4}{3}\gamma_{M}G_{M}(p)^{2}}\right)$
$\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\gamma_{M}G_{M}(p)+\frac{4}{3}\gamma_{M}G_{M}(p)}\right)\leq 2\exp\left(-\frac{nG_{M}(p)}{10\gamma_{M}}\right).$
Lemma 10 (Full Variance Expression).
Assume that the noise $\varepsilon_{i}$ have finite fourth moments. Then, there exists a constant $C>0$ such that
$\mathbb{E}[(\hat{P}_{M}(f)-P_{M}(f))^{2}]$ $\leq 2\frac{\mathbb{E}[\gamma_{M}^{2}\exp(-(X_{i}-x)^{T}M(X_{i}-x))^{2}\{Y_{i}-P_{M}(f)\}^{2}]}{nG_{M}(p)^{2}}$
$\quad\quad+2\frac{\mathbb{E}[\{\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x))Y_{i}\}^{2}]}{nG_{M}(p)^{2}}+C\exp\left(-\frac{nG_{M}(p)}{10\gamma_{M}}\right)$
Proof.
Let $E$ be the event $|\hat{C}_{M}-G_{M}(p)|\leq G_{M}(p)/2$. We compute
$\hat{P}_{M}(f)-P_{M}(f)$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x))Y_{i}}{\hat{C}^{\prime}_{M}}-\frac{G_{M}(fp)}{G_{M}(p)}$
$=\frac{\frac{1}{n}\sum_{i=1}^{n}[\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))\{Y_{i}-P_{M}(f)\}]}{G_{M}(p)}\mathbf{1}_{E}$
$\quad\quad+\frac{G_{M}(p)-\hat{C}^{\prime}_{M}}{G_{M}(p)\hat{C}^{\prime}_{M}}\frac{1}{n}\sum_{i=1}^{n}[\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}]\mathbf{1}_{E}$
$\quad\quad+\left\{\frac{\frac{1}{n}\sum_{i=1}^{n}\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}}{\hat{C}^{\prime}_{M}}-\frac{G_{M}(fp)}{G_{M}(p)}\right\}\mathbf{1}_{E^{c}}.$
Hence,
$\mathbb{E}[(\hat{P}_{M}(f)-P_{M}(f))^{2}]$ $\leq 2\frac{\mathbb{E}[\gamma_{M}^{2}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))^{2}\{Y_{i}-P_{M}(f)\}^{2}]}{nP_{M}(p)^{2}}$
$\quad\quad+2\mathbb{E}\left[\left\{\frac{G_{M}(p)-\hat{C}^{\prime}_{M}}{G_{M}(p)\hat{C}^{\prime}_{M}}\frac{1}{n}\sum_{i=1}^{n}[\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}]\right\}^{2}\mathbf{1}_{E}\right]$
$\quad\quad+2\mathbb{E}\left[\left\{\hat{P}_{M}(f)-P_{M}(f)\right\}^{2}\mathbf{1}_{E^{c}}\right].$
The first term is as desired. Next, by definition of the event, we have
$\mathbb{E}\left[\left\{\frac{G_{M}(p)-\hat{C}^{\prime}_{M}}{G_{M}(p)\hat{C}^{\prime}_{M}}\frac{1}{n}\sum_{i=1}^{n}[\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}]\right\}^{2}\mathbf{1}_{E}\right]\leq\frac{\mathbb{E}[\left\{\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}\right\}^{2}]}{nG_{M}(p)^{2}}.$
For the remaining term, observe that
$\hat{P}_{M}(f)=\sum_{i=1}^{n}w_{i}Y_{i}$
where $w_{i}$ are a convex combination. Hence,
$\mathbb{E}\left[\left\{\hat{P}_{M}(f)-P_{M}(f)\right\}^{2}\mathbf{1}_{E^{c}}\right]\leq\sqrt{\mathbb{E}\left[\left\{2\hat{P}_{M}(f)^{2}+2P_{M}(f)^{2}\right\}^{2}\right]}\mathbb{P}[E^{c}]\leq C\exp\left(-\frac{nG_{M}(p)}{10\gamma_{M}}\right),$
by Lemma 9 and the hypothesis on the moments of $\varepsilon_{i}$ and the boundedness of $f$. ∎
With this out of the way, we bound the two remaining expectations.
Lemma 11 (Normalization Constant).
There exists a constant $C>0$ such that
$\mathbb{E}[\gamma_{M}^{2}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))^{2}\{Y_{i}-P_{M}(f)\}^{2}]$ $\leq C\gamma_{M}G_{2M}(p),$
$\mathbb{E}[\{\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}\}^{2}]$ $\leq C\gamma_{M}G_{2M}(p).$
Proof.
Notice that $\gamma_{M}/\gamma_{4M}=C$ for some constant. Let $\varepsilon\sim\rho$. Focusing on the first expression,
$\mathbb{E}[\gamma_{M}^{2}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))^{2}\{Y_{i}-P_{M}(f)\}^{2}]$
$\quad\quad=C\gamma_{M}\int\gamma_{4M}\exp(-(z-x)^{T}[2M](z-x))p(y)(f(y)+\varepsilon-P_{M}(f))^{2}\,dz\,d\rho(\varepsilon)$
$\quad\quad=C\gamma_{M}\mathbb{E}_{2M}[p(x+Z)\mathbb{E}_{\rho}[(f(x+Z)+\mathcal{E}-P_{M}(f)(x))^{2}]]\leq C^{\prime}\gamma_{M}G_{2M}(p)(x),$
using bounded moments and bounded $f$ assumptions. Similarly,
$\mathbb{E}[\{\gamma_{M}\exp(-(X_{i}-x^{*})^{T}M(X_{i}-x^{*}))Y_{i}\}^{2}]$ $=C\gamma_{M}\mathbb{E}_{2M,\rho}[p(Z)[f(Z)+\mathcal{E}]^{2}]=C\gamma_{M}G_{2M}(p).$
Combining these yields the desired variance bound multiplied by density dependent factors.
Proof of Lemma 2.
By Lemmas 10 and $1$, there exists a constant $C>0$ such that
$\mathbb{E}[(\hat{P}_{M}(f)-P_{M}(f))^{2}]\leq C\frac{\gamma_{M}}{n}\left(\frac{G_{2M}(p)}{G_{M}(p)^{2}}\right)+C\exp\left(-\frac{n}{\gamma_{M}}\left(\frac{G_{M}(p)}{10}\right)\right)\\ =O\left(\frac{\sqrt{\det M}}{n}\right).$
B.1.3 MSE
We compile the results from the previous section to provide guarantees for the MSE and MISE against Gaussians.
Proof of Theorem 1.
Compiling Lemmas 1 and 2,
$\mathbb{E}_{M}[(\hat{P}_{M}(f)-f)^{2}]$
$\quad\quad\leq C\frac{\sqrt{\det M}}{n}\left(\frac{G_{2M}(p)}{G_{M}(p)^{2}}\right)+C\exp\left(-\frac{n}{\sqrt{\det M}}\left(\frac{G_{M}(p)}{10}\right)\right)+C\left(1+\frac{\operatorname{Var}_{M_{t}}(p(Z))}{G_{M_{t}}(p)^{2}}\right)W(\mu_{M})$
$\quad\quad=O\left(\frac{\sqrt{\det M_{t}}}{n}+W(\mu_{t})\right).$
B.1.4 General Kernels
In this section, we discuss which of results above are applicable to kernels of the form $k(M^{1/2}x)$. We start with the following Poincaré inequality. Let $\mu_{k,M}$ denote the probability distribution with density $\propto k(M^{1/2}[x-x^{*}])$.
Lemma 12 (Kernel Poincaré).
The measure $\mu_{k,M}$ has covariance proportionate to $M^{-1}$. If $\mu_{k,I}$ satisfies a Poincaré inequality with constant $1/C_{I}$, then
$\operatorname{Var}_{\mu_{k,M}}(f)\leq\frac{1}{C_{I}}\int\nabla f^{T}M^{-1}\nabla f\,d\mu_{k,M}.$
Proof.
For convenience, we translate $x^{*}=0$ without loss of generality. For the first claim, notice that when $M=I$, the density is symmetric about the origin, hence the covariance is proportionate to the identity. Thus the claim follows as $\mu_{k,M}$ is the pushforward of $\mu_{k,I}$ by $M^{-1/2}$.
Further,
$\operatorname{Var}_{\mu_{k,M}}(f)$ $=\frac{1}{\sqrt{\det{M}}}\operatorname{Var}_{\mu_{k,I}}(f\circ M^{-1/2})\leq\frac{1}{C_{I}}\int\frac{1}{\sqrt{\det M}}\|\nabla f(M^{-1/2}y)\|^{2}d\mu_{k,I}(y).$
An application of the chain rule gives that
$\nabla_{y}f(M^{-1/2}y)=M^{-1/2}\nabla f(M^{-1/2}y).$
Substituting this into the bound above gives
$\operatorname{Var}_{\mu_{k,M}}(f)$ $\leq\frac{1}{C_{I}}\int\frac{1}{\sqrt{\det M}}\|M^{-1/2}\nabla f(M^{-1/2}y)\|^{2}\,\mu_{k,I}(y).$
Changing variable back to $\mu_{k,M}$ yields the desired result. ∎
With this in hand, the following is immediate for $\hat{P}_{M}$ with $k$ generic. The claim for the covariance similarly follows by change of variables, after observing that $\mu_{k,I}$ is rotation invariant, and therefore has isotropic covariance.
Corollary 2 (Kernel MSE).
Take $k$ to be nonnegative, and satisfying a Poincaré inequality for isotropic covariance. Assume that $f$ is bounded and $\varepsilon_{i}$ has finite fourth moments. Then there exists an absolute constant $C$ such that
$\mathbb{E}\left[\int(\hat{P}_{M}(f)-f)^{2}\,d\mu_{k,M}\right]\leq$
$\quad\quad C\frac{\sqrt{\det M}}{n}\left(\frac{\mathbb{E}_{\mu_{k^{2},2M}}[p]}{\mathbb{E}_{\mu_{k,2M}}[p]^{2}}\right)+C\exp\left(-\frac{n}{\sqrt{\det M}}\left(\frac{\mathbb{E}_{\mu_{k,2M}}[p]}{10}\right)\right)+C\left(1+\frac{\operatorname{Var}_{\mu_{k,M}}(p)}{\mathbb{E}_{\mu_{k,M}}[p]^{2}}\right)W(\mu_{k,M}).$
B.2 Recursive Kernel Learning as Variance Minimization
The proof of Proposition 1 is clear from the main text, but we fill in the missing details here.
Proof of Proposition 1.
The only thing left to clarify is the identity
$tA^{-1}/D=\operatorname{argmax}_{X\succeq 0,\operatorname{tr}(AX)\leq t}\log\det X.$
Note that $\log\det X$ is strictly convex, and KKT conditions yield that the minimum is a solution to the Lagrangian system of equations. Evaluating, we have
$\mathcal{L}(X,\lambda)=\log\det X-\lambda(\operatorname{tr}(AX)-t)\Longrightarrow\nabla_{X}\mathcal{L}=X^{-1}-\lambda A=0.$
Hence the optimum is proportionate to $A^{-1}$, and evaluating the constraint yields the result. ∎
B.3 Guarantees
To derive the learning rate for the Local EGOP Learning algorithm, we mostly operate at the level of matrices, studying the recurrence from Equation 7. To supplement this in the noisy manifold setting, we rely on critical geometric relations verified in Appendix B.3.2.
B.3.1 Sample Trimming
A practical concern when implementing the Local EGOP Learning algorithm is that two points $p,p^{\prime}\in\mathcal{M}$ may be such that $p-p^{\prime}\in\mathcal{N}_{p}$, but $f(p)\neq f(p^{\prime})$. Such a function can still be continuous index as the line segment between these points goes beyond the reach of the manifold, and thus outside the data sampling region. A simple fix is to exclude data points outside of a fixed isotopic radius of the point of interest $x^{*},$ thus trimming the population distribution $P.$ The resulting EGOP matrix will then be defined with respect to an extension outside this domain. We adopt such an extension so that Lemma 17, and the secondary conclusions of Lemma 5, that there is an affine space where the gradient is constant and the Hessian is degenerate, hold globally. Alternatively, one can broach this by taking the covariance cap to 0 at a logarithmic rate, reducing the sampling weight of the non-local region at the expense of a small variance inflation. For simplicity, we adopt the former assumptions.
B.3.2 Supporting Arguments for Lemma 4
We try to keep our proofs as elementary as possible. We start with basic properties of our function of interest $f$.
Lemma 13 (Derivatives).
Suppose that $f=f\circ\pi$. Then $\nabla f(x)\in\mathcal{T}_{x}$ and for any $u,v\in\mathcal{N}_{x}$, $u^{T}\nabla^{2}f(x)v=0$.
Proof.
Let $u,v\in\mathcal{N}_{x}$ be arbitrary. Then,
$\nabla f(x)^{T}v$ $=\lim_{t\to 0}[f(x+tv)-f(x)]/t=\lim_{t\to 0}[f(\pi[x+tv])-f(\pi[x])]/t=\lim_{t\to 0}[f(\pi[x])-f(\pi[x])]/t=0,$
as for small $t$ orthogonal perturbations do not change the nearest point on $\mathcal{M}$. As this holds for generic $v$ the first claim follows. To simplify notation, define $\partial_{\eta}h(x):=\lim_{t\to 0}[h(x+t\eta)-h(x)]/t$. We can thus compute
$u^{T}\nabla^{2}f(x)v=\partial_{u}(\nabla f^{T}v)(x)=0$
as $\nabla f^{T}v$ is identically 0 along the path $x+tv$, $t\in[0,\delta]$ for $\delta$ small enough. That is, as $\nabla f$ remains tangent along this path, as verified above, it is always orthogonal to $v$. ∎
In the interest of further studying the structure of $\nabla f$ we verify the following fundamental Geometric facts. We denote by $\pi_{V}$ the projection onto the vector space $V$. Denote by $d(\cdot,\cdot)$ the geodesic distance on $\mathcal{M}$. Let $\kappa=\sup_{p\in\mathcal{M},\|v\|=1}\|S_{v}(p)\|$, $S$ the Weingarten map of $\mathcal{M}$. This constant is automatically finite as $\mathcal{M}$ has nontrivial reach.
Lemma 14 (Orthogonal Curvature).
There exists a universal constant $C$ such that, for any $p,p^{*}\in\mathcal{M}$,
$\|\pi_{\mathcal{N}_{p}}-\pi_{\mathcal{N}_{p^{*}}}\|\leq Cd(p,p^{*}).$
Proof.
Let $\gamma_{t}$, $t\in[0,d(p,p^{*})]$ be a unit speed geodesic connecting $p$ and $p^{*}$. We can compute
$\|\pi_{\mathcal{N}_{p}}-\pi_{\mathcal{N}_{p^{*}}}\|=\left\|\int_{0}^{1}\partial_{t}\pi_{\mathcal{N}_{\gamma_{t}}}\,dt\right\|\leq\int_{0}^{d(p,p^{*})}\|\partial_{t}\pi_{\mathcal{N}_{\gamma_{t}}}\|\,dt.$
Thus, if we can show a uniform bound for $\|\partial_{t}\pi_{\mathcal{N}_{\gamma_{t}}}\|$ then we are done.
Let $v_{1},\dots,v_{D-d}$ be an orthonormal basis of $\mathcal{N}_{x^{*}}$, and form the orthonormal moving frame $v_{1}(t),\dots,v_{D-d}(t)$ along $\gamma_{t}$ by parallel translation. Let
$V(t):=\begin{bmatrix}v_{1}(t)&\dots&v_{D-d}(t)\end{bmatrix}$.
By construction, $\pi_{\mathcal{N}_{\gamma_{t}}}=V(t)V(t)^{T}$, $V(t)$ is orthogonal for all $t$. Differentiating,
$\|\partial_{t}\pi_{\mathcal{N}_{\gamma_{t}}}\|=\|V^{\prime}(t)V(t)^{T}+V(t)V^{\prime}(t)^{T}\|\leq 2\|V^{\prime}(t)\|\|V(t)\|=2\|V^{\prime}(t)\|.$
This reduces the problem to uniformly bounding $\|V^{\prime}(t)\|$. The key relations to prove this are
$V^{\prime}(t)=\begin{bmatrix}v_{1}^{\prime}(t)&\dots&v_{D-d}^{\prime}(t)\end{bmatrix},\quad v_{i}^{\prime}(t)=-S_{v_{i}(t)}(\gamma_{t})\gamma^{\prime}_{t}.$
The first identity is definitional, the second can be found in Do Carmo [2016]. For $\|u\|=1$, we have
$\|V^{\prime}(t)u\|=\|\sum_{i=1}^{D-d}u_{i}v_{i}^{\prime}(t)\|=\|-S_{\sum_{i}u_{i}v_{i}(t)}(\gamma_{t})\gamma_{t}^{\prime}\|\leq\|S_{\sum_{i}u_{i}v_{i}(t)}(\gamma_{t})\gamma_{t}^{\prime}\|\leq\kappa,$
as $\|\sum_{i}u_{i}v_{i}(t)\|\leq\sum_{i}u_{i}^{2}=1$, hence $\|V^{\prime}(t)\|\leq\kappa$ holds uniformly. ∎
Lemma 15 (Projection Normal).
For any $v\in\mathcal{N}_{\pi[x]}$, $D\pi_{x}v=0$.
Proof.
As computed previously,
$D\pi_{x}v=\lim_{t\to 0}[\pi[x+tv]-\pi[x]]/t=0.$
Lemma 16 (Orthogonal Control).
Fix $x^{*}$ in the tubular neighborhood of $\mathcal{M}$. There exists $r>0$ small enough such that, for $\|x-x^{*}\|\leq r$, there exists a constant $C>0$ such that
$\|\pi_{\mathcal{N}_{x^{*}}}-\pi_{\mathcal{N}_{x}}\|\leq C\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|$
Proof.
By Lemma 14, it suffices to show that $d(\pi[x^{*}],\pi[x])=O(\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|).$ By Garcia Trillos [2019], we have that $d(\pi[x^{*}],\pi[x])=O(\|\pi[x^{*}]-\pi[x]\|)$, further reducing the problem. By our locality assumption, we can take convex combinations and remain within the tubular neighborhood, allowing us to decompose
$\|\pi[x^{*}]-\pi[x]\|\leq\left\|\int_{0}^{1}D\pi_{x^{*}+t(x-x^{*})}(x-x^{*})\,dt\right\|\leq\int_{0}^{1}\|D\pi_{x^{*}+t(x-x^{*})}(x-x^{*})\|\,dt.$
By Lemma 18,
$\|D\pi_{x^{*}+t(x-x^{*})}(x-x^{*})\|=\|D\pi_{x^{*}+t(x-x^{*})}(\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}])\|\leq C\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}]\|$
$\quad\quad\leq C\|[\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}][x-x^{*}]\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|\leq Cr\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|,$
as $\pi$ is smooth, and we have restricted ourselves to a bounded domain. In other words,
$\|\pi[x^{*}]-\pi[x]\|-Cr\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|\leq C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|,$
hence, if we can show that
$\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|\leq C^{\prime}\|\pi[x^{*}]-\pi[x]\|,$
then for $r$ small enough the desired inequality will be achieved. Notice that $\pi_{\mathcal{T}_{x}}=I-\pi_{\mathcal{N}_{x}}$, hence by Lemma 14
$\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|=\|\pi_{\mathcal{N}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{N}_{x^{*}}}\|\leq\tilde{C}d(\pi[x^{*}],\pi[x^{*}+t(x-x^{*})])\leq C^{\prime}\|\pi[x^{*}]-\pi[x]\|,$
completing the proof. ∎
We apply this to study $\nabla f$.
Lemma 17 (Gradient Rotation).
Restricting ourselves to a region small enough that Lemma 16 is valid, we have
$\|\pi_{\mathcal{N}_{x^{*}}}\nabla f(x)\|=O(\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|).$
Proof.
By Lemmas 13 and 16,
$\|\pi_{\mathcal{N}_{x^{*}}}\nabla f(x)\|=\|[\pi_{\mathcal{N}_{x^{*}}}-\pi_{\mathcal{N}_{x}}]\nabla f(x)\|\leq C\|\pi_{\mathcal{N}_{x^{*}}}-\pi_{\mathcal{N}_{x}}\|\leq C^{\prime}\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|.$
B.3.3 Supporting Arguments for Lemma 5
Lemma 18.
Let $x\in\mathcal{M}$ and $\eta\in\mathcal{N}_{x}$. Then $D\pi_{x+\eta}=(I-S_{x}(\eta))^{-1}\pi_{\mathcal{T}_{x}}$.
Proof.
Clearly, for any normal vector $v\in\mathcal{N}_{x}$, $D\pi_{x+\eta}v=0$, as projection is invariant to normal displacement. Thus it suffices to verify the formula for any tangent vector. Our proof has two primary parts. First, we pass to normal coordinates, equating the tubular neighborhood about $\mathcal{M}$ with the normal bundle, and compute the differential of the exponential map. From this parameterization, we can then analyze the projection via the chain rule.
Starting from the normal bundle $N\mathcal{M}$, for $(x,v)\in N_{x}\mathcal{M}$, the exponential map $E(x,v)=x+v$ is a diffeomorphism to the tube. Let $\gamma(t)=(u(t),\xi(t))\in N\mathcal{M}$, $\gamma^{\prime}(t)\in TN\mathcal{M}$. We compute $\xi^{\prime}(0)=-S_{x}(v)u^{\prime}(0)+\nabla_{u^{\prime}(0)}^{\perp}\xi(0)$, hence
$DE_{(x,v)}\gamma^{\prime}(0)=u+[-S_{x}(v)u+\nabla_{u^{\prime}(0)}^{\perp}\xi(0)]=(I-S_{x}(v))u+\nabla_{u^{\prime}(0)}^{\perp}\xi(0).$
Composing with the projection map, $\pi\circ E(x,v)=x$, hence
$D\pi_{x+v}DE_{(x,v)}\gamma^{\prime}(0)=u^{\prime}(0)\quad\Longleftrightarrow\quad D\pi_{x+v}\pi_{\mathcal{T}_{x}}[(I-S_{x}(v)u]=u^{\prime}(0),$
and inverting yields the claim, as $D\pi_{x+v}$ is a linear operator. ∎
Proof of Lemma 5.
The first claim follows immediately from Lemma 18 and the chain rule, indeed,
$\nabla f(x+v)=\nabla g(\pi(x+v))=(I-S_{x}(v))^{-1}\nabla g(x).$
Let $u=(I-S_{x}(v))^{-1}\nabla g(x)$. Applying linearity of the shape operator, observe that $(I-S_{x}(v+\eta))u=(I-S_{x}(v))u-S_{x}(\eta)(u)=\nabla g(x)-S_{x}(\eta)u$. Hence, if $S_{x}(\eta)u=0$ then $u=\nabla f(x+v+\eta)$. Thus, the second claim follows if we can show that there is a D2dD-2d dimensional subspace $\operatorname{Null}$ that annihilates the solution vector $u$. For $w\in\mathcal{T}_{x}$,
$\langle S_{x}(\eta)u,w\rangle=\langle\mathrm{I\!I}(u,w),\eta\rangle$
for $\mathrm{I\!I}$ the second fundamental form. This is a bilinear map, and the map defined by fixing the $u$ argument, $\phi_{u}(w):=\mathrm{I\!I}(u,w)$ is a linear map from $\mathcal{T}_{x}\to\mathcal{N}_{x}$, and thus has image no larger than $d$. Its complement $\phi_{u}(\mathcal{T}_{x})^{\perp}$ has dimension at least $\operatorname{dim}(\mathcal{N}_{x})-d=D-2d$, as desired.
This proves the invariance of the gradient along a space of dimension at least D2dD-2d, which immediately implies $\nabla^{2}f(x+v+\eta)z=\partial_{z}\nabla f(x+v)=0.$
B.3.4 Taylor Expansion
We first provide the generic Taylor expansion, which we then go on to refine for the noisy manifold setting.
Proof of Lemma 3.
We start by Taylor expanding the gradient. For a $D$ tuple $\xi$, as $\partial_{\xi}\partial_{i}f=\partial_{i}\partial_{\xi}f$, for $f$ sufficiently smooth we can express
$\partial_{i}f(x)=\sum_{|\xi|\leq q}\frac{1}{q!}\partial_{i}[\partial_{\xi}f](x)\prod_{q_{i}\in\xi}x_{i}^{q_{i}}+O(\|x\|^{q+1}).$
In particular, we have
$\partial_{i}f(x)=\partial_{i}f(0)+\sum_{j}\partial_{i}\partial_{j}f(0)x_{j}+\frac{1}{2}\sum_{j,k}\partial_{i}\partial_{j}\partial_{k}f(0)x_{j}x_{k}+O(\|x\|^{3}).$
Working at the level of the gradient, this yields
$\nabla f(x)=\nabla f(0)+\nabla^{2}f(0)x+T(x,x)+O(\|X\|^{3})$
where $T$ is the bilinear for
$T[x,y]]_{i}=\frac{1}{2}\sum_{j,k}\partial_{i}\partial_{j}\partial_{k}f(0)x_{j}y_{k}.$
We use the first three terms of this Taylor expansion as a starting point. Set $g:=\nabla f(0),H:=\nabla^{2}f(0)$.
The remainder of the Taylor expansion can be exactly expressed as
$\tilde{R}(x)=\frac{1}{6}\int_{0}^{1}(1-s)^{2}\nabla^{4}f(sx)[x,x,x]\,ds,\quad[\nabla^{4}f(sx)[x,x,x]]_{i}:=\sum_{j,k,\ell}\partial_{i}\partial_{j}\partial_{k}\partial_{\ell}f(sx)x_{i}x_{j}x_{k}.$
Hence, under our uniform boundedness assumption, we have
$\|T[x,x]\|\leq C\|x\|^{2},\quad\|\tilde{R}(x)\|\leq C\|x\|^{3}$
for some constant $C$. Throughout the argument, for convenience we will absorb constants into $C$ if necessary. Hence,
$\mathcal{L}(\mu)=\int\nabla f\nabla f^{T}\,d\mu=gg^{T}+H\Sigma H+gT(\Sigma)^{T}+T(\Sigma)g^{T}+R(\Sigma),$
for
$T(\Sigma)_{i}$ $:=\mathbb{E}_{\mu}[T(x,x)_{i}]=\frac{1}{2}\sum_{j,k}\partial_{i}\partial_{j}\partial_{k}f(0)\Sigma_{jk}$
$R(\Sigma)$ $:=\mathbb{E}_{\mu}[T[Z,Z]T[Z,Z]^{T}+HZ\tilde{R}(Z)^{T}+\tilde{R}(Z)(HZ)^{T}+\tilde{R}(Z)\tilde{R}(Z)^{T}].$
That $T$ is continuous and linear in $\Sigma$ is immediate. For the remainder, we have
$\|R(\Sigma)\|$ $\leq\mathbb{E}_{\mu}[\|T[Z,Z]T[Z,Z]^{T}\|+\|HZ\tilde{R}(Z)^{T}\|+\|\tilde{R}(Z)(HZ)^{T}\|+\|\tilde{R}(Z)\tilde{R}(Z)^{T}\|]$
$\leq C\mathbb{E}[\|Z\|^{4}+\|Z\|^{6}]\leq C\mathbb{E}_{\mu}[\operatorname{tr}(\Sigma^{2})+\operatorname{tr}(\Sigma)^{2}]$
where we apply Isserlis’s Theorem Isserlis [1918] to relate moments of the norm to the covariance matrix, and our capped covariance assumption is used to compress all higher order terms to quadratic order. Our proof is completed upon observing
$\operatorname{tr}(\Sigma^{2})=\|\Sigma\|_{F}^{2},\quad\operatorname{tr}(\Sigma)^{2}=\left(\sum_{i}\lambda_{i}\right)^{2}\leq D\sum_{i}\lambda_{i}^{2}=D\|\Sigma\|_{F}^{2},$
and in finite dimensions the frobenius and operator norms are equivalent. ∎
Proof of Lemma 4.
That $\operatorname{rank}(H)\leq 2d$ is immediate from Lemma 13. Indeed, the $\mathcal{N}\times\mathcal{N}$ block of $H$ is completely 0, hence at most $d$ rows and $d$ columns of $H$ in the $\mathcal{T}\oplus\mathcal{N}$ basis have nonzero entries. Let $B_{r}$ be
We now apply Lemmas 13 and 17 to compute
$\|\pi_{\mathcal{N}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|=\left\|\int\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\,d\mu\right\|\leq\int\|\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\|\,d\mu=\int\|[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu$
$\quad\quad\leq C\int\|\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}\|^{2}\,d\mu\leq C\int\|\pi_{\mathcal{T}}x\|^{2}\,d\mu\leq C[\operatorname{tr}([\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}}]^{2})+\operatorname{tr}(\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}})^{2}]\leq C\|\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}}\|_{F}^{2},$
where the final bounds from Isserlis’s Theorem are similar to Lemma 3. Similarly,
$\|\pi_{\mathcal{T}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|\leq\int\|\pi_{\mathcal{T}}\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu\leq C\int\|\pi_{\mathcal{T}}x\|\,d\mu\leq C\sqrt{\int\|\pi_{\mathcal{T}}x\|^{2}\,d\mu}=C\|\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}}\|_{F}.$
As the operator and Frobenius norms are equivalent the bounds are as desired. ∎
Proof of Corollary 1.
The basic claim is immediate from Lemma 3, and that the terms stemming from higher order derivatives are uniformly bounded follows from the smoothness assumptions on $f$. Further, the nullspace of the Hessian and the 0 order expansion being the gradient $g=\nabla f(x^{*})$ follow immediately from Lemma 5. ∎
B.3.5 Matrix Recurrence
We begin our study of the recurrence in Equation 7 by introducing a reduced form. We analyze this scheme, and then apply it to Local EGOP Learning, leveraging the results of Appendix B.3.2. We divide our focus into two distinct settings, the first mirroring Equation 9 by introducing a constant shift, and the second mirroring Equation 8 a homogeneous equation.
Constant Shift
Let $J$ be an involution, $J=J^{T}=J^{-1}.$ We introduce the recurrence
$B_{i+1}=t_{i+1}J(G+\beta B_{i}+(1-\beta)B_{i-1}+E_{i})^{-1}J$
where $E_{i}$ is an error term, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $t_{1}=\alpha^{2}$, $B_{1}=t_{1}J(G+B_{0})^{-1}J$, $t_{i+1}=t_{i}r^{2}$, $r<1$.
Lemma 19 (Perturbed Shift).
Suppose there exists a uniform constant $C$ such that $\|E_{i}\|\leq C\alpha$ for all $i$. Then, for $\alpha$ small enough,
$\|B_{i+1}\|=\Theta(t_{i}).$
Proof.
That $\alpha$ must be small is not necessary, but this will be a typical setting and will help expedite the proof. Let $\alpha<1$ be small enough that $\|B_{i}\|\leq\lambda_{\min}(G)/4$, $i=1,2$, and $C\alpha\leq\lambda_{\min}(G)/4$. We prove by strong induction that $\|B_{i+1}\|\leq t_{i+1}\frac{2}{\lambda_{\min}(G)}$. The base case is already proven, so we move on to the inductive step. We can simply compute
$G+\beta B_{i}+(1-\beta)B_{i-1}+E_{i}\succeq G/2.$
As $J$ is an orthogonal matrix, we have $\|B_{i+1}\|\leq 2t_{i+1}/\lambda_{\min}(G)$ and the claim follows.
With the decay of $B_{i}$ being settled, it follows that $B_{i}=t_{i+1}JG^{-1}J+O(t_{i+1}^{2}+\alpha)$, hence, normalizing by $t_{i+1}$, we have exact convergence to $JG^{-1}J$ in the noiseless setting. ∎
Note that the geometric decay and many of the assumptions were not completely necessary for this proof, however we introduce them now as they will greatly facilitate the following arguments.
Homogeneous
We move on to the recurrence
$A_{i+1}=t_{i+1}J(\beta A_{i}+(1-\beta)A_{i-1}+E_{i})^{-1}J$
where $E_{i}$ is an error term, $J$ is an involution, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $t_{1}=\alpha^{2}$, $B_{1}=t_{1}J(G+B_{0})^{-1}J$, $t_{i+1}=t_{i}r^{2}$, $r<1$.
We start by reducing this iteration scheme to an autonomous equation.
Proposition 2 (Autonomous Recurrence).
Define $c(r,\beta):=\frac{1}{\beta/r+(1-\beta)/r^{2}}$ and let $X_{i}:=A_{i}/\sqrt{c(r,\beta)t_{i}}$, $R_{i}:=\sqrt{c(r,\beta)}E_{i}/\sqrt{t_{i+1}}$, and $\eta=\frac{\beta/r}{c(r,\beta)}$. Then $X_{i}$ satisfies the autonomous recurrence
$X_{i+1}=J(\eta X_{i}+(1-\eta)X_{i-1}+R_{i})^{-1}J$
Proof.
We work this out by simple algebra. First, as an intermediate step, introduce $C_{i}:=A_{i}/\sqrt{t_{i}}$, $R_{i}^{\prime}:=E_{i}/\sqrt{t_{i+1}}$. Then we can write
$A_{i+1}=t_{i+1}J(\beta A_{i}+(1-\beta)A_{i-1}+E_{i})^{-1}J\Longleftrightarrow\frac{A_{i+1}}{\sqrt{t_{i+1}}}=J\left(\beta\frac{A_{i}}{\sqrt{t_{i+1}}}+(1-\beta)\frac{A_{i-1}}{\sqrt{t_{i+1}}}+\frac{E_{i}}{\sqrt{t_{i+1}}}\right)^{-1}J$
$\quad\quad\Longleftrightarrow C_{i+1}=J([\beta/r]C_{i}+[(1-\beta)/r^{2}]C_{i-1}+R_{i}^{\prime})^{-1}J\Longleftrightarrow C_{i+1}=c(r,\beta)J(\eta C_{i}+(1-\eta)C_{i-1}+R_{i}^{\prime}c(r,\beta))^{-1}J$
$\quad\quad\Longleftrightarrow\frac{C_{i+1}}{\sqrt{c(r,\beta)}}=J\left(\eta\frac{C_{i}}{\sqrt{c(r,\beta)}}+(1-\eta)\frac{C_{i}}{\sqrt{c(r,\beta)}}+R_{i}^{\prime}\sqrt{c(r,\beta)}\right)J\Longleftrightarrow X_{i+1}=J(\eta X_{i}+(1-\eta)X_{i-1}+R_{i})^{-1}J.$
Thus to study $A_{i}$ it suffices to study $X_{i}$, then multiply by the corresponding factors. Our analysis proceeds as follows. First, leveraging the Thompson metric [Thompson, 1963], we show that, for summable error $R_{i}$, the spectrum of $X_{i}$ stays bounded away from 0 and $\infty$. Then, utilizing this bound, we go on to show that $X_{i}\to I$ up to some error controlled by $\alpha$. Define the Thompson metric,
$d_{T}(A,B):=\log\max\{M(A/B),M(B/A)\},\quad M(A/B):=\inf\{\gamma>0:A\preceq\gamma B\}.$
Lemma 20 (Spectral Band).
Suppose that $\sum_{i}\|R_{i}\|\leq C\alpha$ and $\max\{d_{T}(X_{0},I),d_{T}(X_{1},I)\}=L_{0}$. Then, for $\alpha$ sufficiently small,
$e^{-L_{0}}-O(\alpha)]I\preceq X_{i}\preceq[e^{L_{0}}+O(\alpha)]I$
for all $i$.
Proof.
Notice that
$(1-\xi_{i})[\eta X_{i}+(1-\eta)X_{i-1}]\preceq\eta X_{i}+(1-\eta)X_{i-1}+R_{i}\preceq(1+\xi_{i})[\eta X_{i}+(1-\eta)X_{i-1}]$
for $\xi_{k}:=\|[\eta X_{i}+(1-\eta)X_{i-1}]^{-1/2}R_{i}[\eta X_{i}+(1-\eta)X_{i-1}]^{-1/2}\|.$ It follows that
$d_{T}([\eta X_{i}+(1-\eta)X_{i-1}]+R_{i},[\eta X_{i}+(1-\eta)X_{i-1}])\leq\log\frac{1+\xi_{i}}{1-\xi_{i}}$
Hence, if $e^{-L_{i}}I\preceq X_{i},X_{i-1}\preceq e^{L_{i}}I$, it follows that $\xi_{i}\leq e^{L_{i}}\|R_{i}\|$ and
$L_{i+1}\leq L_{i}+\log\frac{1+\|R_{i}\|e^{L_{i}}}{1-\|R_{i}\|e^{L_{i}}}\Longleftrightarrow e^{-L_{i+1}}\geq e^{-L_{i}}\left(\frac{e^{-L_{i}}+\|R_{i}\|}{e^{-L_{i}}-\|R_{i}\|}\right)\Longrightarrow e^{-L_{i+1}}\geq e^{-L_{i}}-2\|R_{i}\|.$
Thus,
$e^{-L_{i}}\geq e^{-L_{0}}-2\sum_{i=1}^{\infty}\|R_{i}\|\geq e^{-L_{0}}/2$
choosing $\alpha$ sufficiently small. ∎
Lemma 21 (Non-Expansive).
Let $X_{i}$, $X_{i}^{\prime}$ be autonomous homogeneous recurrences of the Proposition 2 form, with common conjugation matrix $J$. Then
$d_{T}(X_{i+1},X_{i+1}^{\prime})\leq\max\{d_{T}(X_{i},X_{i}^{\prime}),d_{T}(X_{i-1},X_{i-1}^{\prime})\}+O(\|R_{i}\|+\|R_{i}^{\prime}\|)$
Proof.
Set $M_{i}=\eta X_{i}+(1-\eta)X_{i-1},M_{i}^{\prime}=\eta X_{i}^{\prime}+(1-\eta)X_{i-1}^{\prime}$. Using the invariance of the Thompson metric to conjugation by orthogonal matrices and inversion, we can compute
$d_{T}(X_{i+1},X_{i+1}^{\prime})$ $=d_{T}(\eta X_{i}+(1-\eta)X_{i-1}+R_{i},\eta X_{i}^{\prime}+(1-\eta)X_{i-1}^{\prime}+R_{i}^{\prime})$
$\leq d_{T}(M_{i},M_{i}^{\prime})+d_{T}(M_{i},M_{i}+R_{i})+d_{T}(M_{i}^{\prime},M_{i}^{\prime}+R_{i}^{\prime}).$
Checking each term,
$d_{T}(M_{i},M_{i}^{\prime})\leq\max\{d_{T}(X_{i},X_{i}^{\prime}),d_{T}(X_{i-1},X_{i-1}^{\prime})\}$
as averaging in the Thompson metric is non-expansive.
$d_{T}(A,A+A)=\|\log A^{-1/2}(A+E)A^{-1/2}\|=\|\log(I+A^{-1/2}EA^{-1/2})\|=O(\|E\|)$
for $A$ with non-vanishing spectrum. Lemma 20 verifies that the sequences $X_{i},X_{i}^{\prime}$ do not degenerate for $\alpha$ sufficiently small, thus showing the claim. ∎
Lemma 22 (Definite Hessian).
Let $J=I$. Then for $\alpha$ sufficiently small, $X_{i}\to I.$
Proof.
Define $T(x,y)=((\eta x+(1-\eta)y)^{-1},x)$, hence $T(X_{i},X_{i-1})=(X_{i+1},X_{i})$. Define
$V(x,y)=\max\{d_{T}(x,I),d_{T}(y,I)\}=\max\{|\log\lambda_{\max}(x)|,|\log\lambda_{\max}(y)|,|\log\lambda_{\min}(x)|,|\log\lambda_{\min}(y)|\}.$
Notice that
$\min\{\lambda_{\min}(x),\lambda_{\min}(y)\}\leq\lambda_{\min}(\eta x+(1-\eta)y)^{-1}\leq\lambda_{\max}(\eta x+(1-\eta)y)^{-1}\leq\max\{\lambda_{\max}(x),\lambda_{\max}(y)\}$
hence $V(T(X_{i},X_{i-1}))\leq V(X_{i},X_{i-1})$. Thus the noiseless recurrence is autonomous, and the discrete Lasalle invariance principle [Mei and Bullo, 2017, Theorem 2] yields the existence of a set $E_{c}$, $T(E_{c})\subseteq E_{c}$ such that for all $(X,Y)\in E_{c}$, $V(E_{c})=c$ and $\lim_{i\to\infty}\inf_{E\in E_{c}}\|(X_{i},X_{i-1})-E\|\to 0$. To verify that $X_{i}\to I$ in the noiseless case, it suffices to show $E_{0}=\{(I,I)\}$ is the only invariant set with constant value function.
We argue by contradiction. Let $E_{c}$ be such a limit set, $c>0$. Take $(X_{1},X_{0})\in E_{c}$, and consider $(X_{3},X_{2})=T^{2}(X_{1},X_{0})$. By the condition on the value function $V$, either the largest eigenvalue of these matrices is $e^{c}$ or the smallest is $e^{-c}$. Without loss of generality, assume that $X_{3}$ achieves the eigenvalue $e^{c}$.
For this to occur, it must be that both $X_{2},X_{1}$ have eigenvalue $e^{-c}$ with a common eigenvector $v$. As the sequence is rational in $X_{2},X_{1}$, $v$ will be an eigenvector for all subsequent matrices $X_{i}$. Notice $\log\lambda_{v}(X_{4})=-\log(\eta e^{c}+(1-\eta)e^{-c})\in(-c,c)$, hence $\sup_{i>0}|\log\lambda_{v}(X_{3+i})|<c$, and this eigenspace can never again be extremal. Indeed, for $\lambda_{v}(X_{3+i})=e^{\pm c}$, we require $\lambda_{v}(X_{2+i})=\lambda_{v}(X_{1+i})=e^{\mp c}$, thus any extremal eigenspace for $T^{2}(X_{1},X_{0})$ will decay upon further iterations, and non-extremal eigenspaces cannot become extremal. In particular, $V(T^{k}(X_{1},X_{0}))$ must decrease as $k\to\infty$, contradicting that $E_{c}$ is invariant with fixed value function $V(E)=c$ for all $E\in E_{c}$.
In the noisy setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can apply [Mischaikow et al., 1995, Theorem 1.8] to verify that the perturbed recurrence converges to the same limit sets as the autonomous recurrence, i.e. $I$ is the unique limit point.
Lemma 23 (Generic J).
Suppose that $J$ has signature $(k,j)$. Then, for $\alpha$ sufficiently small, $X_{i}\to L\in\{X:(XJ)^{2}=I\}$. Further, $L$ has at most $2\min\{k,j\}$ eigenvalues not equal to 1.
Proof.
We update the recurrence in Lemma 22 by $J$ congruence. $V$ is still monotone in the autonomous noiseless recurrence as $J$ is orthogonal. We begin by investigating the structure of any invariant set $E_{c}$.
As in the previous proof, if we initialize with a pair in $E_{c}$, then for any $i\geq 2$, and all extremal eigenpairs $(\lambda_{\ell},v_{\ell})$, it must be that for both j=i1,i2j=i-1,i-2, the matrices $X_{j}$ have matching eigenpairs $(1/\lambda_{\ell},Jv_{\ell})$. Repeating this argument for $X_{i+1}$, for any extremal eigenpairs $(\nu_{\ell},u_{\ell})$ it follows that $X_{i},X_{i-1}$ have matching eigenpairs $(1/\nu_{\ell},Ju_{\ell})$. Combining this information, for each extremal eigenpair $(\nu_{\ell},u_{\ell})$ of $X_{i+1}$, $(1/\nu_{\ell},Ju_{\ell})$ must be an eigenpair of both $X_{i}$ and $X_{i-1}$. As $(\lambda_{\ell},v_{\ell})=(1/\nu_{\ell},Ju_{\ell})$ is an extremal eiegenpair of $X_{i}$, it must also be that $(\nu_{\ell},u_{\ell})$ is an eigenpair of $X_{i-1}$. Hence the matrix $X_{i-1}$ has both $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ as extremal eigenpairs, and for them to pass forward in the recurrence they must be shared by $X_{i-1}.$ Thus we see that for all $m\geq i$, $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ are fixed eigenvectors of $X_{m}$.
Thus we have proven that any limiting set in $E_{c}$ has fixed, cyclical extremal eigenpairs. Suppose that $X_{i}$ has $\kappa$ such limiting maximal eigenvalues. Let $\operatorname{Trunc}_{\kappa}(X)$ be the map that identifies $X$ with the $(D-\kappa)\times(D-\kappa)$ matrix formed by removing the principal $\kappa$ components. Thus, $\operatorname{Trunc}_{\kappa}T(X_{i+1},X_{i})=T(\operatorname{Trunc}_{\kappa}(X_{i},X_{i-1}))+o(1)$, and the sequence $\operatorname{Trunk}_{\kappa}(X_{i},X_{i-1})$ is asymptotically autonomous and monotone in the value function $V$. Hence [Mischaikow et al., 1995, Theorem 1.8] yields that the truncated sequence has asymptotically cyclical extremal eigenvectors.
Inductively applying this argument verifies that $X_{i}\to L$ for a fixed matrix $L$ in the noiseless case. By definition
$L=\lim_{i\to\infty}X_{i+1}=\lim_{i\to\infty}J(\eta X_{i+1}+(1-\eta)X_{i})^{-1}J=JL^{-1}J,$
hence $L\in\{X:(LJ)^{2}=I\}.$ In the noisy setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can apply [Mischaikow et al., 1995, Theorem 1.8] to verify that the perturbed recurrence converges to the same limit sets as the autonomous recurrence.
We now study the structure of such $L$. As we demonstrated, every eigenpair $(\lambda,v)$ of $L$ must have matching eigenpair $(1/\lambda,Jv)$. Let $v_{1},\dots,v_{\kappa},Jv_{1},\dots,Jv_{\kappa},u_{1},\dots,u_{D-2\kappa}$, where $Ju_{i}=\pm u_{i}$ are invariant under $J$. For each eigenvector $u_{i}$, $\lambda=1$ is forced, thus distinct eigenvalues can only come from the 2-cycles. Displaying $J$ in this basis, we have
$J=\begin{bmatrix}0&I&0\\ I&0&0\\ 0&0&D\end{bmatrix}$
The first $2\kappa\times 2\kappa$ block has signature $(\kappa,\kappa)$, yielding the desired claim. ∎
Lemma 24 (Second Order Convergence).
Let $r\in(0,1)$, $\|R_{i}\|\leq C\alpha r^{i}$, $X_{0},X_{1}$ commute with $J$. Then, for any $\varepsilon>0$, there exists $\delta>0$ such that $X_{i}\to L_{\alpha}$, $\|L_{\alpha}-I\|<\varepsilon$.
Proof.
That $X_{i}$ has a limit $L_{\alpha}$ for $\alpha$ small follows immediately from Lemma 23. Thus it suffices to verify that $\lim_{i\to\infty}X_{i}=I$ for the unperturbed sequence. Indeed, if for $i$ large enough, $\|X_{i}-I\|<\varepsilon,\|X_{i-1}-I\|<\varepsilon$, because $X_{i}$ is polynomial in $X_{0},X_{1},R_{j}$, $j=1,\dots,i-1$, it follows that it is smooth for $\alpha$ small, thus incurring additional $O(\alpha)$ error. In combination with Lemma 20, this verifies the claim.
Without perturbations, the whole sequence commutes, and
$J(\eta X_{i}+(1-\eta)X_{i-1})^{-1}J=(\eta X_{i}+(1-\eta)X_{i-1})^{-1}J^{2}=(\eta X_{i}+(1-\eta)X_{i-1})^{-1}$
hence $J$ factors out, and the convergence is a consequence of Lemma 22. ∎
B.3.6 Full Rank Hessian
Our main proof strategy is to reduce the Local EGOP iterations to the form of Proposition 2, so that we can guarantee stable first and second order convergence rates by Lemmas 19 and 24. A straightforward change of variables yields the following. Let $H$ have diagonalization $H=U\Lambda U^{T}$.
Proposition 3 (H Reduction).
Let $J=\operatorname{sign}(\Lambda)$ be the diagonal signature matrix and $W=U|\Lambda|^{1/2}$.
Define the transformation
$X_{i}:=W^{T}\Sigma_{i}W,\qquad\tilde{g}:=JW^{-1}g,$
$\tilde{T}(X):=JW^{-1}T(W^{-T}XW^{-1})W^{-T}J,\qquad\tilde{R}(X):=JW^{-1}R(W^{-T}XW^{-1})W^{-T}J.$
Then the recurrence for $X_{i}$ depends only on $J$:
$X_{i+1}=t_{i+1}J\Big($ $\tilde{g}\tilde{g}^{T}+\beta X_{i}+(1-\beta)X_{i-1}$
$+\beta[\tilde{g}\tilde{T}(X_{i})^{T}+\tilde{T}(X_{i})\tilde{g}^{T}+\tilde{R}(X_{i})]$
$+(1-\beta)[\tilde{g}\tilde{T}(X_{i-1})^{T}+\tilde{T}(X_{i-1})\tilde{g}^{T}+\tilde{R}(X_{i-1})]\Big)^{-1}J$
Thus we can reduce our recurrence to that studied in Section B.3.5.
Lemma 25 (Schur Complement).
[…]
$\alpha_{i}:=u^{T}\Sigma_{i}u,\quad\alpha_{i}^{\prime}:=u^{T}H\Sigma_{i}Hu,\quad\eta^{2}:=\|g\|^{2},\quad b_{i}:=\bar{Q}\Sigma_{i}u,\quad b_{i}^{\prime}:=\bar{Q}H\Sigma_{i}Hu,$
$C_{i}:=\bar{Q}\Sigma_{i}\bar{Q},\quad C_{i}^{\prime}:=\bar{Q}H\Sigma_{i}H\bar{Q},\quad R_{b,i}:=\eta\bar{Q}T(\Sigma_{i})+\bar{Q}R(\Sigma_{i})u,$
$\bar{R}_{\alpha,i}:=\beta[2\eta u^{T}T(\Sigma_{i})+u^{T}R(\Sigma_{i})u]+(1-\beta)[2\eta u^{T}T(\Sigma_{i-1})+u^{T}R(\Sigma_{i-1})u],$
$\bar{b}_{i}=\beta[b_{i}^{\prime}+R_{b,i}]+(1-\beta)[b_{i-1}^{\prime}+R_{b,i-1}],\qquad\bar{R}_{C,i}:=\beta\bar{Q}R(\Sigma_{i})\bar{Q}+(1-\beta)\bar{Q}R(\Sigma_{i-1})\bar{Q},$
$\bar{S}_{i}:=\beta H_{\bar{Q}}C_{i}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{i-1}H_{\bar{Q}}+\bar{R}_{C,i}-\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}.$
Then, in the orthonormal basis $\{u\}\oplus\mathrm{range}(\bar{Q})$,
$\begin{bmatrix}\alpha_{i+1}&b_{i+1}^{T}\\ b_{i+1}&C_{i+1}\end{bmatrix}=t_{i+1}$
$\begin{bmatrix}\displaystyle\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}+\left(\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\right)^{2}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}&\displaystyle-\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\\ \displaystyle-\bar{S}_{i}^{-1}\bar{b}_{i}\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}&\displaystyle\Big(\beta C_{i}^{\prime}+(1-\beta)C_{i-1}^{\prime}+\bar{R}_{C,i}-\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big)^{-1}\!\end{bmatrix}$.
Proof.
This is the Schur complement with pivot on the top-left scalar applied to $M_{i}:=[\beta\mathcal{L}(\mu_{i})+(1-\beta)\mathcal{L}(\mu_{i-1})]^{-1}$. ∎
Corollary 3 (Base Case).
Suppose that $\Sigma_{0}=\alpha I$ and $t_{1}=\alpha^{2}$. Then, in the $(u,\bar{Q})$ basis,
$\Sigma_{1}=\begin{bmatrix}\frac{\alpha^{2}}{\eta^{2}}+O(\alpha^{3})&O(\alpha^{2})\\ O(\alpha^{2})&\alpha(\bar{Q}H^{2}\bar{Q})^{-1}+O(\alpha^{2})\end{bmatrix},\quad\Sigma_{2}=\begin{bmatrix}\frac{\alpha^{2}}{\eta^{2}}+O(\alpha^{3})&O(\alpha^{2})\\ O(\alpha^{2})&\alpha H_{\bar{Q}}^{-1}H^{2}H_{\bar{Q}}^{-1}+O(\alpha^{2})\end{bmatrix}$
Proof.
This can be directly computed from the formula in Lemma 25. As a subtle detail, at the first step, the covariance is isotropic, hence in the $\bar{Q}\times\bar{Q}$ block the dominant terms is $\alpha(\bar{Q}H(I)H\bar{Q})^{-1}=\alpha(\bar{Q}H^{2}\bar{Q})^{-1}$, however, after this step the gradient element of the covariance is a higher order residual, being of the Lemma 19 constant shift form, and thus subsequently $\bar{Q}H\Sigma_{1}H\bar{Q}=(\bar{Q}H\bar{Q})\Sigma_{1}(\bar{Q}\Sigma_{1}\bar{Q})+O(\alpha^{2})=:H_{\bar{Q}}\Sigma_{1}H_{\bar{Q}}+O(\alpha^{2}).$
Proof of Theorem 1.
We seek to inductively apply Lemmas 19 and 24 to the $u\times u$ and $\bar{Q}\times\bar{Q}$ blocks respectively, necessitating a careful tabulation of the residuals and off-diagonals. Our inductive hypothesis is that
$\alpha_{i}=\Theta(t_{i}),\quad b_{i}=O(\sqrt{t_{i}}),\quad C_{i}=\Theta(\sqrt{t_{i}}),\quad\bar{R}_{C,i}=O(t_{i}),\quad\bar{R}_{b,i}=O(\sqrt{t_{i}}),\quad\bar{R}_{\alpha,i}=O(\sqrt{t_{i}}).$
We argue by strong induction, with the base case being covered in Corollary 3. We argue one block at a time. First note that $\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}=O(\sqrt{t_{i}})$, hence
$\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}+\left(\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\right)^{2}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}=\frac{1}{\eta^{2}}+O(\sqrt{t_{i}})\Longrightarrow\alpha_{i+1}=\frac{t_{i+1}}{\eta^{2}}+O\left(t_{i+1}^{3/2}\right).$
Thus, taking $\alpha$ sufficiently small, we get a uniform $O(t_{i})$ bound by Lemma 19 and Proposition 3.
For the $\bar{Q}\times\bar{Q}$ block,
$C_{i+1}=t_{i+1}\Big(\beta H_{\bar{Q}}C_{i}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{i-1}H_{\bar{Q}}+\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{i}+(1-\beta)\Sigma_{i-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big]\Big)^{-1}.$
As the rescaled residual
$\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{i}+(1-\beta)\Sigma_{i-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big]/\sqrt{t_{i}}\leq C\sqrt{t_{i}}=C\alpha r^{i}$
is uniformly summable in $\alpha$, we can apply Propositions 2 and 3, and Lemma 23 to verify $C_{i+1}=\Theta(\sqrt{t_{i}})$.
Moving on to the off-diagonal, $\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}=O(1)$, thus upon multiplying by $t_{i+1}$, this is $O(t_{i})$, whereas the subsequent remainder terms $R_{b,i+1}=\Theta(\|\Sigma_{i+1}\|)=O(\sqrt{t_{i+1}})$. Thus, the contribution is negligible at the following iteration, and the uniform rate follows.
We can immediately plug these rates into our risk bounds to yield $\sqrt{\det\Sigma_{i}}=\Theta(t_{i}^{(D+1)/4})$, $W(\mu_{t})=O(t_{i})$, thus yielding the desired rate. ∎
B.3.7 Noisy Manifold
We start off by iterating Corollary 1 as described in the text.
Lemma 26 (Weak Dependence).
Suppose that for all $v\in\operatorname{Null}$, $\eta\in\operatorname{Null}^{\perp}$, $\eta^{T}\Sigma v=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})$. Then, for $\Sigma_{\operatorname{Null}^{\perp}}:=\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}^{\perp}}$, $\Sigma_{\operatorname{Null}}:=\pi_{\operatorname{Null}}\Sigma\pi_{\operatorname{Null}}$ there exists $H_{\Sigma_{\operatorname{Null}}},T_{\Sigma_{\operatorname{Null}}}$ such that
$\mathcal{L}(\mu)=gg^{T}+H_{\Sigma_{\operatorname{Null}}}\Sigma_{\operatorname{Null}^{\perp}}H_{\Sigma_{\operatorname{Null}}}+gT_{\Sigma_{\operatorname{Null}}}(\Sigma_{\operatorname{Null}^{\perp}})^{T}+T_{\Sigma_{\operatorname{Null}}}(\Sigma_{\operatorname{Null}^{\perp}})g^{T}+R_{\Sigma_{\operatorname{Null}}}(\Sigma_{\operatorname{Null}^{\perp}}),$
$\|T_{\Sigma_{\operatorname{Null}}}(\Sigma_{\operatorname{Null}^{\perp}})\|=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|),\ \|R_{\Sigma_{\operatorname{Null}}}(\Sigma_{\operatorname{Null}^{\perp}})\|=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})$.
Proof.
Applying our Gaussian ansatz,
$\Sigma_{v}:=\operatorname{Cov}_{N(x^{*},\Sigma)}(X|\pi_{\operatorname{Null}}[X-x^{*}]=v)=\Sigma_{\operatorname{Null}^{\perp}}-[\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}}]\Sigma_{\operatorname{Null}}^{-1}[\pi_{\operatorname{Null}}\Sigma\pi_{\operatorname{Null}^{\perp}}]=\Sigma_{\operatorname{Null}^{\perp}}+O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2}).$
Thus, by Corollary 1,
$\mathcal{L}(\mu_{v})$ $=gg^{T}+H_{v}\Sigma_{v}H_{v}+gT_{v}(\Sigma_{v})^{T}+T_{v}(\Sigma_{v})g^{T}+R_{v}(\Sigma_{v})$
$=gg^{T}+H_{v}\Sigma_{\operatorname{Null}^{\perp}}H_{v}+gT_{v}(\Sigma_{\operatorname{Null}^{\perp}})^{T}+T_{v}(\Sigma_{\operatorname{Null}^{\perp}})g^{T}+R_{v}(\Sigma_{\operatorname{Null}^{\perp}})+O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2}),$
and this latter term can be absorbed into the remainder. Now, integrating out $v$ the relevant operators are
$H_{\Sigma_{\operatorname{Null}}}=\int H_{\pi_{\operatorname{Null}[x-x^{*}]}}\,d\mu(x),\quad T_{\Sigma_{\operatorname{Null}}}=\int T_{\pi_{\operatorname{Null}[x-x^{*}]}}\,d\mu(x),\quad R_{\Sigma_{\operatorname{Null}}}=\int R_{\pi_{\operatorname{Null}[x-x^{*}]}}\,d\mu(x).$
While this matrix depends on the $\operatorname{Null}$ component of $\Sigma$, we will develop a recurrence where this is consistent across iterations. We now expand Lemma 25 in the low-rank Hessian setting.
Lemma 27 (Low-rank Schur).
For $\Sigma=\alpha I$ operators are defined as in Lemma 25, otherwise their equivalents as introduced in Lemma 26. Decomposing in the $(\bar{Q}_{\operatorname{Null}^{\perp}},\bar{Q}_{\operatorname{Null}}):=(\operatorname{col}(\overline{Q})\cap\operatorname{Null}^{\perp},\operatorname{col}(\bar{Q})\cap\operatorname{Null})$ basis, let
$\tilde{R}_{C,i}:=\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T},\quad\bar{C}_{i}:=\beta C_{i}^{\prime}+(1-\beta)C_{i}^{\prime},\quad A_{i}=\bar{Q}_{\operatorname{Null}^{\perp}}\bar{C}_{i}\bar{Q}_{\operatorname{Null}^{\perp}}+\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}},$
$W_{i}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}-\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}\bar{[}Q_{\operatorname{Null}^{\perp}}\bar{C}_{i}\bar{Q}_{\operatorname{Null}^{\perp}}]^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}.$
$C_{i+1}=t_{i+1}\begin{bmatrix}A_{i}^{-1}+A_{i}^{-1}[\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}]W_{i}^{-1}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}]A_{i}^{-1}&-A_{i}^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}W_{i}^{-1}\\ -W_{i}^{-1}Q_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}A_{i}^{-1}&W_{i}^{-1}\end{bmatrix}$
Proof.
Recall that $C_{i+1}:=t_{i+1}(\bar{C}_{i}+\tilde{R}_{C,i})^{-1}$, hence this is the block Schur complement in the specified basis pivoted on the $\bar{Q}_{\operatorname{Null}^{\perp}}\times\bar{Q}_{\operatorname{Null}^{\perp}}$ block. ∎
Corollary 4 ($\bar{Q}$ block).
Suppose that $\Sigma_{0}=\alpha I$ and $t_{1}=\alpha^{2}$. Then, in the $(\bar{Q}_{\operatorname{Null}^{\perp}},\bar{Q}_{\operatorname{Null}})$ basis
$C_{1}=\begin{bmatrix}\alpha(\bar{Q}_{\operatorname{Null}^{\perp}}H^{2}\bar{Q}_{\operatorname{Null}^{\perp}})^{-1}+O(\alpha^{2})&O(\alpha)\\ O(\alpha)&\alpha^{2}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,0}\bar{Q}_{\operatorname{Null}}]^{-1}+O(\alpha)\end{bmatrix}$,
$C_{2}=\begin{bmatrix}\alpha H_{\bar{Q}_{\operatorname{Null}^{\perp}}}^{-1}H^{2}H_{\bar{Q}_{\operatorname{Null}^{\perp}}}^{-1}+O(\alpha^{2})&O(\alpha)\\ O(\alpha)&\alpha^{2}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,1}\bar{Q}_{\operatorname{Null}}]^{-1}+O(\alpha)\end{bmatrix}$.
The matrix $\alpha^{2}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}]^{-1}=\Theta(1)$, $i=0,1$.
Proof.
We can once again check this via direct computation utilizing Lemma 27. As $\bar{C}_{0}$ comprises terms linear in $\Sigma_{0}$, it is $\Theta(\alpha)$, meanwhile the remainder terms $\tilde{R}_{C,0}=O(\alpha^{2})$. Hence $A_{0}=\bar{Q}_{\operatorname{Null}^{\perp}}\bar{C}_{0}\bar{Q}_{\operatorname{Null}^{\perp}}+O(\alpha^{2})$, $W_{i}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,0}\bar{Q}_{\operatorname{Null}}+O(\alpha^{3})$, and the remaining approximations directly follow. As in the proof of Lemma 3, at the first step, the covariance is isotropic, hence in the $\bar{Q}_{\operatorname{Null}^{\perp}}\times\bar{Q}_{\operatorname{Null}^{\perp}}$ block the dominant terms is $\alpha(\bar{Q}_{\operatorname{Null}^{\perp}}H(I)H\bar{Q}_{\operatorname{Null}^{\perp}})^{-1}=\alpha(\bar{Q}_{\operatorname{Null}^{\perp}}H^{2}\bar{Q}_{\operatorname{Null}^{\perp}})^{-1}$, however, after this step the gradient element of the covariance is a higher order residual, being of the Lemma 19 constant shift form, and thus subsequently $\bar{Q}_{\operatorname{Null}^{\perp}}H\Sigma_{1}H\bar{Q}_{\operatorname{Null}^{\perp}}=(\bar{Q}_{\operatorname{Null}^{\perp}}H\bar{Q}_{\operatorname{Null}^{\perp}})\Sigma_{1}(\bar{Q}_{\operatorname{Null}^{\perp}}H\bar{Q}_{\operatorname{Null}^{\perp}})+O(\alpha^{2})=:H_{\bar{Q}_{\operatorname{Null}^{\perp}}}\Sigma_{1}H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(\alpha^{2}).$
Corollary 5 (Noisy Manifold Base Case).
Suppose that $X_{0}=\alpha I$ and $t_{1}=\alpha^{2}$. In the basis $(g,\overline{Q}_{\operatorname{Null}^{\perp}},\overline{Q}_{\operatorname{Null}}),$ we have
$\Sigma_{1}=\begin{bmatrix}\frac{\alpha^{2}}{\eta^{2}}+O(\alpha^{3})&O(\alpha^{2})&O(\alpha)\\ O(\alpha^{2})&\alpha(\bar{Q}_{\operatorname{Null}^{\perp}}H^{2}\bar{Q}_{\operatorname{Null}^{\perp}})^{-1}+O(\alpha^{2})&O(\alpha)\\ O(\alpha)&O(\alpha)&\Theta(1)\end{bmatrix}$,
$\Sigma_{2}=\begin{bmatrix}\frac{\alpha^{2}}{\eta^{2}}+O(\alpha^{3})&O(\alpha^{2})&O(\alpha)\\ O(\alpha^{2})&\alpha H_{\bar{Q}_{\operatorname{Null}^{\perp}}}^{-1}H^{2}H_{\bar{Q}_{\operatorname{Null}^{\perp}}}^{-1}+O(\alpha^{2})&O(\alpha)\\ O(\alpha)&O(\alpha)&\Theta(1)\end{bmatrix}$.
Proof.
Corollary 4 accounts for the bottom right $2\times 2$ block, what remains is to verify the top and left rows. Specifically, we must verify how
$C_{1}(b_{0}+R_{b,0}):=\alpha^{2}\left(\alpha H_{\bar{Q}}^{2}+\bar{R}_{C,0}-\frac{1}{\eta^{2}+\alpha_{0}+\bar{R}_{\alpha,0}}\bar{b}_{0}\bar{b}_{0}^{T}\right)^{-1}(b_{0}+R_{b,0})$
decomposes across $\bar{Q}_{\operatorname{Null}^{\perp}},\bar{Q}_{\operatorname{Null}}$. Applying Corollary 4, we can express this block-wise as
$\begin{bmatrix}O(\alpha^{2})\\ O(\alpha)\end{bmatrix}=\begin{bmatrix}O(\alpha)&O(\alpha)\\ O(\alpha)&O(1)\end{bmatrix}\begin{bmatrix}O(\alpha)\\ O(\alpha)\end{bmatrix}$
yielding the desired rates. The proof for $\Sigma_{2}$ is essentially the same. ∎
Define $\mu(\zeta)=N(x^{*},[\Sigma-\Sigma_{\operatorname{Null}^{\perp}}]+\zeta\bar{Q}_{\operatorname{Null}})$, i.e. the distribution with $\operatorname{Null}\times\operatorname{Null}$ component fixed to $\zeta$.
Lemma 28 (Filtered $\operatorname{Null}$).
Assume that $D\geq 2d$ and $\operatorname{rank}(H)=2d$. Let
$\Sigma_{i+1}=t_{i+1}[\beta\mathcal{L}(\mu_{i}(\zeta))+(1-\beta)\mathcal{L}(\mu_{i-1}(\zeta)]^{-1}$
$\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, for $v\in\operatorname{Null}$, $v^{T}\Sigma_{i}v=\Theta(1)$, and for $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\theta(\sqrt{t})$.
Proof.
Applying Corollary 1, we can express
$\mathcal{L}(\mu)=\int\mathcal{L}(\mu_{v})\,d\nu(v)=\int H_{v}\Sigma_{v}H_{v}+gT_{v}(\Sigma_{v})^{T}+T(\Sigma_{v})^{T}+R_{v}(\Sigma_{v})\,d\nu(v),$
where $\nu$ is the marginal density $\nu(v)=p_{\mu}(\pi_{\operatorname{Null}}[X-x^{*}]=v)$.
We can apply Corollary 5 to verify that the desired anisotropy is present starting from an isotropic initialization. We first study the modified recurrence where the metric $M$ is in terms of $\mathcal{L}^{\prime}(\mu)$ rather than the EGOP. We proceed by induction to verify that these rates are maintained. Our inductive hypothesis is that
$\alpha_{i}=\Theta(t_{i}),\quad b_{i}=O(\sqrt{t_{i}}),\quad C_{i}=\Theta(\sqrt{t_{i}}),\quad\bar{R}_{C,i}=O(t_{i}),$
$\bar{R}_{b,i}=O(\sqrt{t_{i}}),\quad\bar{R}_{\alpha,i}=O(\sqrt{t_{i}}),\quad\|\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma\bar{Q}_{\operatorname{Null}}\|=O(\sqrt{t_{i}}).$
The weak $\operatorname{Null}$, $\operatorname{Null}^{\perp}$ corelation allows us to apply Lemma 26 to reduce the recurrence to the marginal form
$\mathcal{L}(\mu_{\zeta})=gg^{T}+H_{\zeta}\Sigma_{\operatorname{Null}^{\perp}}H_{\zeta}+gT_{\zeta}(\Sigma_{\operatorname{Null}^{\perp}})^{T}+T_{\zeta}(\Sigma_{\operatorname{Null}^{\perp}})g^{T}+R_{\zeta}(\Sigma_{\operatorname{Null}^{\perp}}).$
By assumption $H_{0}=H$ is maximal rank, hence for $\zeta$ sufficiently small this property is maintained for $H_{\zeta}$, and we restrict our analysis to this neighborhood. The verification for the $g\times g$ block is exactly the same as verified in Theorem 1 as it is of constant shift form (see Lemma 19), thus we move on to the remaining terms. Explicit computations are similar to those previously presented in Corollaries 4 and 5. First, we immediately have that
$A_{i}=H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta C_{i}+(1-\beta)C_{i-1}]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(t_{i}),\quad W_{i}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}+O(t_{i}^{3/2}).$
Thus for the middle block, applying Lemma 27 yields
$\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}\bar{Q}_{\operatorname{Null}^{\perp}}=t_{i+1}\{H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta(\bar{Q}_{\operatorname{Null}^{\perp}}C_{i}\bar{Q}_{\operatorname{Null}^{\perp}})+(1-\beta)(\bar{Q}_{\operatorname{Null}^{\perp}}C_{i-1}\bar{Q}_{\operatorname{Null}^{\perp}})]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(t_{i})\}^{-1}.$
by the reductions from Propostions 2 and 3, we can reduce this to an autonomous recurrence in $Q_{\operatorname{Null}^{\perp}}C_{i}\bar{Q}_{\operatorname{Null}^{\perp}}$ that converges at the desired rate by Lemma 23. See the the proof of Theorem 1 for more details.
Applying Lemma 27 again yields
$\bar{Q}_{\operatorname{Null}}C_{i+1}\bar{Q}_{\operatorname{Null}}=t_{i+1}W_{i}^{-1}=t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{t_{i}}).$
As $\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})=O(t_{i})$, it follows that $t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})^{-1}=\Omega(1)$, and thus it is of constant order by our covariance cap. Thus this rate is an immediate consequence of the second order anisotropy, and the orthogonal component and other remainders only contribute negligibly, justifying the uniform rate. The remaining terms in the Lemma 27 block decomposition are
$\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma_{i+1}\bar{Q}_{\operatorname{Null}}=\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}\bar{Q}_{\operatorname{Null}}=-t_{i+1}A_{i}^{-1}(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})W_{i}^{-1},\quad b_{i+1}=u^{T}\Sigma_{i+1}\bar{Q}=C_{i+1}(b_{i}+R_{b,i}),$
for $u=g/\|g\|$ the normalized gradient vector. In the first term, the product $(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})W_{i}^{-1}=\Theta(1)$, hence the rate is dominated by the second order anisotropy $t_{i+1}A_{i}^{-1}=\Theta(\sqrt{t_{i}})$. For the final vector, $\|b_{i}+R_{b,i}\|=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|)$, hence we can combine this with our previous analysis to yield
$\bar{Q}_{\operatorname{Null}}C_{i+1}(b_{i}+R_{b,i})=\Theta(1)O(\sqrt{t_{i}}))=O(\sqrt{t_{i+1}}),\quad\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}(b_{i}+R_{b,i})=\Theta(t_{i+1})O(\sqrt{t_{i}})=O(t_{i+1}).$
Note that the scalar inflation to go from $t_{i}$ to $t_{i+1}$ does not aggregate across iterations as the initial bound was driven by the second order anisotropy term, to which the remainders only contribute negligibly.
This verifies the claim for the recurrence in $\mathcal{L}(\mu_{\zeta})$. Our goal now is to lift this analysis to $\mathcal{L}(\mu)$ with covariance cap $\zeta$. The distinction between these recurrences is that the first a priori filters the $\operatorname{Null}$ direction out of the dynamics, whereas the latter trims any eigenvalues above a set threshold $\zeta$. These recurrences coincide if the $\operatorname{Null}\times\operatorname{Null}$ component never falls below the threshold $\zeta$, thus we seek to identify such a threshold.
Proof of Theorem 3.
Lemma 28 verifies the specified dynamics, particularly the $\operatorname{Null}\times\operatorname{Null}$ component of $\Sigma_{i,\zeta}$ remains $\Theta(1)$, and thus there is some floor $\gamma(\zeta)$ which it never drops below. We now will argue that $\gamma(\zeta)>0$ as $\zeta\to 0$, hence there exists a threshold $\zeta^{*}$ such that the filtered and cap dynamics coincide.
As $\zeta\to 0$, $H_{\zeta}\to H$, as the distribution concentrates about $x^{*}$, and in particular, for $\zeta$ small enough, the signature of $H_{\zeta}$ is the same as that of $H$. This follows from the maximal rank assumption on $H$. By Propositions 2 and 3, upon transforming $C_{i,\zeta}\mapsto\frac{W_{\zeta}^{T}C_{i,\zeta}W_{\zeta}}{\sqrt{c(r,\beta)t_{i}}}=:X_{i,\zeta}$, $W_{\zeta}:=U_{\zeta}|\Lambda_{\zeta}|^{1/2}$, for the diagonalization $H_{\zeta}=U_{\zeta}\Lambda_{\zeta}U_{\zeta}^{T}$, these recurrences are reduced to the same canonical form, with seeds depending smoothly on $\zeta$. Applying Lemma 21, it follows that $X_{i,\zeta}=X_{i,0}+O(\zeta+\alpha)$.
Inverting the transformation, this yields $C_{i,\zeta}=C_{i,0}+O(\sqrt{t_{i}}(\delta+\alpha)).$ Returning to the proof of Lemma 28,
$\Sigma_{i,\operatorname{Null},\zeta}=t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i,\zeta}(\Sigma_{i,\operatorname{Null}^{\perp},\zeta})\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{t_{i}})=\Sigma_{i,\operatorname{Null}^{\perp},0}+O(\sqrt{t_{i}}(\zeta+\alpha)),$
hence for $\zeta,\alpha$ sufficiently small the noise floor is non-vanishing as desired.
Thus the rates of Lemma 28 is achieved. Plugging these rates into our risk bounds yields $\sqrt{\det\Sigma_{i}}=\Theta(t_{i}^{(2d+1)/4})$, $W(\mu_{t})=O(t_{i})$, thus yielding the desired rate.
Appendix C Extended Discussion
The most obvious problem left open in our approach is the question of accurate EGOP estimation. One path to achieve this is to show that function gradients can also be estimated accurately up to the prescribed error rates in the domains of interest. At initialization, this is of course reasonable, as we focus on a small isotropic region for which the analysis is standard, and an estimator can achieve a fixed level of bias at an $O(n^{-1/2})$ rate. However, in the iteration that immediately follows this can no longer be expected. As a canonical setting, consider the second order anisotropy $g^{T}\Sigma_{i}g=\Theta(t_{i}),v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$, $v\perp g$. What must be verified is how the elongation impacts pointwise gradient estimates.
When regressing on such an oblique region, the estimated gradient corresponds to the average of the gradients on the ellipsoid. Thus, even if the anisotropy induces bias, it is reasonable to expect that the AGOP, the aggregate of the resulting estimated gradient outerproducts, is very similar to the EGOP. Indeed, we anticipate that the EGOP off-gradient components are on the order of $\Theta(\sqrt{t_{i}})$, the same magnitude as the anticipated bias correction. The supervised noisy manifold hypothesis does not fundamentally change this calculation, as the additional newly introduced non-decaying components of $\Sigma_{i}$ correspond exactly to the directions where the gradient remains constant, and thus present a benign factor in gradient estimation.